### v1

# DAFE: LLM-Based Evaluation Through Dynamic Arbitration for Free-Form Question-Answering

Abstract
Evaluating Large Language Models (LLMs) free-form generated responses remains a challenge due to their diverse and open-ended nature. Traditional supervised signal-based automatic metrics fail to capture semantic equivalence or handle the variability of open-ended responses, while human evaluation, though reliable, is resource-intensive. Leveraging LLMs as evaluators offers a promising alternative due to their strong language understanding and instruction-following capabilities. Taking advantage of these capabilities, we propose the Dynamic Arbitration Framework for Evaluation (DAFE), which employs two primary LLM-as-judges and engages a third arbitrator only in cases of disagreements. This selective arbitration prioritizes evaluation reliability while reducing unnecessary computational demands compared to conventional majority voting. DAFE utilizes task-specific reference answers with dynamic arbitration to enhance judgment accuracy, resulting in significant improvements in evaluation metrics such as Macro F1 and Cohen’s Kappa. Through experiments, including a comprehensive human evaluation, we demonstrate DAFE’s ability to provide consistent, scalable, and resource-efficient assessments, establishing it as a robust framework for evaluating free-form model outputs.
1 Introduction
The rapid advancements in Large Language Models (LLMs) have propelled the field of natural language processing forward, yet their evaluation remains a challenge (Laskar et al., 2024). In particular, free-form model responses are difficult to evaluate because their correctness depends on understanding the broader context and underlying meaning (Si et al., 2021). Many benchmarks, such as MMLU (Hendrycks et al., 2021), often simplify evaluation by focusing on structured formats (e.g., multiple-choice questions) (Chen et al., 2024). Although effective for certain tasks, such methods rely on log probabilities assigned to predefined options, where the model selects the most likely answer, limiting the range of capabilities that can be assessed (Thakur et al., 2024). This structured approach fails to accommodate the complexity of free-form responses, where multiple valid answers exist (Chang et al., 2024). The rigid, predefined options in such evaluations not only limit the scope of assessment but also overlook the diversity of potential correct responses in free-form tasks (Li et al., 2023; Zhang et al., 2024).
Automatic metrics including lexical matching, n-gram, and neural-based have been widely adopted as scalable solutions for the evaluation of free-form model outputs. Lexical matching methods such as Exact Match (EM) evaluate model predictions by assessing strict lexical alignment between generated outputs and reference answers. However, EM fails to account for semantically equivalent variations in phrasing. For instance, despite their equivalence, EM treats “nuclear weapon” and “atomic bomb” as incorrect. Similarly, n-gram-based metrics (Papineni et al., 2002; Lin, 2004) primarily assess surface-level similarity and often fail to capture semantic equivalence, particularly when lexical or structural diversity conveys the same underlying meaning (Zhu et al., 2023; Chen et al., 2021; Zhang et al., 2020). Neural-based metrics like BERTScore (Zhang et al., 2020) address such limitations by leveraging contextual embeddings to evaluate semantic similarity. However, BERTScore depends on reference quality (Liu et al., 2024) and struggles with domain adaptation and length variations (Zhu et al., 2023). Furthermore, continuous score provider metrics are difficult to interpret (Xu et al., 2023). The limitations in automatic metrics become particularly evident when evaluating instruction-tuned chat models (Doostmohammadi et al., 2024), which tend to produce verbose and diverse responses (Saito et al., 2023; Wang et al., 2024b).
Contrary to automatic metrics, human evaluation provides a more transparent assessment (Chiang and Lee, 2023). However, despite being the “gold standard”, human evaluation is not without its limitations. LLMs’ growing complexity and scale have made recruiting and coordinating multiple human raters increasingly resource-intensive and time-consuming (Mañas et al., 2024). Furthermore, the reliability of human evaluation is additionally challenged by variations in rater expertise and inherent subjectivity that affect reproducibility (Clark et al., 2021; Chiang and Lee, 2023).
Recently, a paradigm shift has emerged where LLMs are utilized to judge the candidate model generations for given tasks (Zheng et al., 2024). This model-based method leverages the instruction-following capabilities of LLMs through evaluation prompts or, in some cases, fine-tuned versions of LLMs that are specifically optimized for evaluation. In this new line of work, research primarily focuses on pairwise comparison (Zheng et al., 2024; Wang et al., 2023; Vu et al., 2024), such as instructing an LLM to judge “which assistant response is better”, and single-answer scoring (Verga et al., 2024) like evaluating summarization task based on predefined criteria (e.g., likability, relevance, etc.) (Chiang and Lee, 2023; Hu et al., 2024; Liu et al., 2023; Chan et al., 2024; Chu et al., 2024).
Inspired by a recent study on self-correction where external feedback helps models identify and correct their mistakes (Gou et al., 2024a), we propose to guide LLM-as-a-judge with human-annotated task-specific reference answers in order to explore the potential of LLMs as an alternative to lexical matching (e.g., EM), neural-based (e.g., BERTScore), and human evaluation for automatic evaluation of free-form model responses. Unlike traditional metrics, an LLM judge can leverage its language understanding and instruction-following capabilities to recognize the correctness of open-ended generations.
We propose the Dynamic Arbitration Framework for Evaluation (DAFE), which employs LLM judges to evaluate free-form model responses. Using a single LLM as a judge, while simple, often leads to inconsistent evaluations, undermining trust in the results. On the other hand, the common practice of using large, universally capable models such as GPT-4 as evaluators makes the evaluation process both slow and costly (Jung et al., 2024; Adlakha et al., 2024; Verga et al., 2024), further limiting its broader applicability. Relying on multiple judges for every evaluation, though more reliable, exacerbates these computational challenges, making such approaches impractical at scale. DAFE offers a middle ground between these approaches by utilizing two complementary primary judges to perform the initial assessment. Only when these judges disagree, is a third independent arbitrator engaged to resolve the conflict. This selective arbitration ensures evaluation reliability and fairness while reducing computational overhead. Our experiments reveal that DAFE achieves significant improvements in metrics such as Macro F1 and Cohen’s kappa. Our key contributions include: a detailed analysis of limitations in conventional metrics for free-form QA, an evaluation of LLM judges with insights into their strengths and errors, a comprehensive human evaluation for benchmarking, and the introduction of DAFE—a scalable framework that improves reliability while minimizing the need for additional evaluators through selective arbitration.
Figure 1: Our proposed Dynamic Arbitration Framework for Evaluation (DAFE). Two primary judges, $J_{1}$ and $J_{2}$, first provide verdicts $V_{i_{1}}$ and $V_{i_{2}}$ for an instance $i$. If agree, that consensus $V_{i}$ is the final decision $D_{i}$. If disagree, a tiebreaker model $J_{t}$ independently produces a verdict $V_{t}$. The final decision $D_{i}$ is then determined via majority voting among $\{V_{i_{1}},V_{i_{2}},V_{t}\}$.
2 Methodology
This section briefly describes the key components of our proposed framework.
2.1 Candidate LLMs
A candidate LLM $\mathcal{C}_{\text{llm}}$ generates output $\bar{y}$ for the given input $x$. We first utilized candidate LLMs to obtain outputs for the given free-form question-answering tasks.
2.2 LLMs-as-a-Judge
A judge $\mathcal{J}_{\text{llm}}$ LLM delivers evaluation or verdict $V$ on candidate LLMs $\mathcal{C}_{\text{llm}}$ outputs $\bar{y}$. The $\mathcal{J}_{\text{llm}}$ evaluates output when prompted with $x$ (i.e., $x\rightarrow\mathcal{A}_{\text{llm}}$) and $\bar{y}$. We utilized the reference answer $r$ and prompted $P$ the $\mathcal{J}_{\text{llm}}$ as:
$P=\{x,\bar{y},r\}$
Utilizing $P$, $\mathcal{J}_{\text{llm}}$ performs the evaluation and delivers a decision as $V=J(P)$. The structure of this $V$ depends on the instructions provided in $P$. For instance, if a binary $V$ is required, $J$ assesses whether $\bar{y}$ is aligned with $r$ given the context $x$ and returns True if $\bar{y}$ is deemed correct, or False if it is not. The evaluation $P$ may vary from zero-shot, where $\mathcal{J}_{\text{llm}}$ receives no prior examples, to few-shot, which includes several related examples, or a chain of thought, encouraging $\mathcal{J}_{\text{llm}}$ to reason stepwise through the problem.
2.3 Dynamic Arbitration Framework for Evaluation (DAFE)
In traditional human evaluation settings, when two annotators disagree on a judgment, a third expert is often called upon to resolve the dispute. Drawing inspiration from this efficient human arbitration practice, we propose the Dynamic Arbitration Framework for Evaluation (DAFE). Rather than immediately employing a large powerful or a closed-source LLMs-as-a-judge, DAFE adopts a cost-efficient approach by beginning with two complementary open-source models as primary judges based on their past performance (Kenton et al., 2024). When these judges reach a consensus, no further evaluation is needed. Only in cases of disagreement is the more powerful LLM engaged as an arbitrator, whose decision then creates a majority verdict. This dynamic approach maintains evaluation quality while minimizing reliance on expensive models. The method also accounts for varying skill levels across different LLMs and tasks (Liang et al., 2024; Sun et al., 2024).
Formally, let $V_{i_{1}}$ and $V_{i_{2}}$ denote the verdicts from the two primary judges for the $i$-th evaluation instance. We define the agreement status $A_{i}$ as:
$A_{i}=\begin{cases}1&\text{if }V_{i_{1}}=V_{i_{2}},\\[6.0pt] 0&\text{otherwise}.\end{cases}$
If $A_{i}=1$, the final decision $D_{i}$ is simply $V_{i}$, the agreed-upon verdict of the primary judges. If $A_{i}=0$, a tiebreaker model provides an additional verdict $V_{t}$. The final decision $D_{i}$ is then obtained via majority voting among $\{V_{i_{1}},V_{i_{2}},V_{t}\}$. Formally:
$D_{i}=\begin{cases}V_{i}&\text{if }A_{i}=1,\\[6.0pt] \text{majority}(\{V_{i_{1}},V_{i_{2}},V_{t}\})&\text{if }A_{i}=0.\end{cases}$
The majority operation selects the verdict that appears at least twice among $\{V_{i_{1}},V_{i_{2}},V_{t}\}$. Since there are three votes, at least two must coincide for a majority.
3 Experiments
We utilize the following settings to examine the performance and reliability of individual LLM judges and DAFE.
3.1 Models
We select open and closed-source instruct models to serve as candidates and judges in our experiment. These include DeepSeek-V3 (DeepSeek-AI et al., 2025), Llama-3.1 70B [^1] https://huggingface.co/meta-llama/Meta-Llama-3.1-70B-Instruct (Meta AI, 2024), GPT-3.5-turbo (Brown et al., 2020), Mistral 7B [^2] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3 (Jiang et al., 2023), and Mixtral 8x7B [^3] https://huggingface.co/mistralai/Mixtral-8x7B-Instructv0.1 (Jiang et al., 2024). We also utilize GPT-4o (OpenAI et al., 2023) and DeepSeek-R1 (DeepSeek-AI et al., 2025) in our ablation experiments. To ensure the reproducibility of our experiments, we set the temperature to 0 for all models under study, as the performance of LLM-based evaluators has been shown to drop when temperature increases (Hada et al., 2024). For our proposed DAFE method, we utilized Mistral 7B and Llama 3.1 70B as primary judges with GPT-3.5-turbo as the tiebreaker. In addition, we experiment with other models as tiebreakers in our ablation experiments. In the rest of the paper, we refer both candidate and judge LLMs as: DeepSeek, Llama, GPT, Mistral, and Mixtral.
3.2 Datasets
We focus on free-form question-answering (QA) since it has widespread practical applications and the critical importance of truthfulness in this domain (Gou et al., 2024a; Evans et al., 2021). In our experiment, we utilize five free-form QA datasets: AmbigQA (Min et al., 2020), FreshQA (Vu et al., 2023), HotpotQA (Yang et al., 2018), Natural Questions (Kwiatkowski et al., 2019), and TriviaQA (Joshi et al., 2017). See Appendix A for details.
3.3 Prompts
We designed generalized (i.e., with minimum instructions) zero-shot prompts with role-playing (Kong et al., 2024) for both candidates and judges. Initially, we prompt candidate LLMs to elicit outputs for the given random samples associated with each dataset.
To evaluate the outputs of candidate LLMs, we prompt judge LLMs for binary verdicts (i.e., True or False) using $P=\{x,\bar{y},r\}$ and instructed to provide a brief explanation for their verdicts (see Appendix D for examples). Binary verdicts explicitly differentiate between correct and incorrect answers, minimize subjective interpretations, and simplify the evaluation process, thus facilitating automatic evaluation. In addition to three key prompt components (i.e., $x$, $\bar{y}$, $r$), we define the role of the judge LLMs as “You are a helpful assistant acting as an impartial judge.” to mitigate biases in judgments (Zheng et al., 2024). We chose not to use few-shot or chain-of-thought prompting strategies to keep the solution robust to a variety of tasks. Previous studies have also shown that in-context examples do not significantly improve the performance of model-based evaluators (Hada et al., 2024; Min et al., 2022).
3.4 Baselines
We establish the following baselines.
Exact Match (EM):
For our selected datasets and also free-form QA tasks, EM serves as a standard lexical matching metric to evaluate candidate LLM performance (Izacard and Grave, 2021; Lewis et al., 2020; Gou et al., 2024b). Due to the verbose nature of LLM-generated responses, we adapt EM to classify an answer as correct if any golden answer $r_{i}\in R$ appears within the generated response $\bar{y}$ (i.e., $r_{i}\subseteq\bar{y}$), rather than requiring complete strict string equality (i.e., $\bar{y}=r_{i}$).
BERTScore:
We use BERTScore (Zhang et al., 2020) which measures similarity by comparing contextualized word embeddings derived from a pre-trained BERT model. This enables the evaluation to focus on semantic correctness rather than exact lexical matches. As BERTScore is based on continuous values between -1 and 1, we set a threshold of $\tau=0.5$ to convert continuous similarity scores into binary 0 and 1. The purpose of this conversion is to allow direct comparison with other evaluation methods. For our implementation, we use the microsoft/deberta-xlarge-mnli [^4] https://huggingface.co/microsoft/deberta-xlarge-mnli model (He et al., 2021).
G-Eval:
In addition to automatic metrics, we also utilize G-Eval (Liu et al., 2023), a reference-free framework that uses GPT-4 to assess the quality of the generated text. In this setting, we modify the evaluation prompt by excluding the reference answer $r$ and directly prompted the evaluator model as $P=\{x,\bar{y}\}$ along with instructions.
Human Evaluation:
It remains the gold standard for assessing the outputs of candidate LLMs. We recruit three graduate students from our academic network, all specialized in natural language processing, to serve as annotators. We provide the input given to the candidate LLMs, reference answers, and candidate LLMs responses. This format, while similar, is distinct from the judge LLMs prompts which additionally require formatted decisions. We anonymize the origin of model responses to reduce potential bias linked to model familiarity or reputation. The annotators were asked to score the candidate LLMs outputs on a binary scale: ‘1’ for ‘True’ and ‘0’ for ‘False’ based on alignment with the reference answer and contextual relevance. For inter-rater reliability, we compute Fleiss’ Kappa ($\kappa$(Fleiss and Cohen, 1973) and percent agreement. See Appendix B for details.
4 Results
Figure 2 illustrates the raw performance of Llama obtained through various evaluators. Unlike lexical matching and neural-based metrics, each LLM-as-a-judge shows overall performance close to the human majority. The proposed DAFE method consistently achieves comparable or slightly better alignment with the human majority. Conventional metrics such as EM severely underestimate the candidate LLMs’ performance. Contrarily, BERTScore tends to overestimate the performance except in some cases such as when evaluating Llama on AmbigQA and NQ-Open (see Table 6 in Appendix C for additional results).
Figure 2: Raw accuracy of candidate Llama across free-form QA tasks using automatic metrics and model-based evaluation. The Human Majority (HM) serves as the ground truth for all evaluators.
4.1 Alignment with human evaluation
We calculate Cohen’s kappa (McHugh, 2012) to find the agreement between each evaluator and the human majority to obtain instance-level comparison. Overall, DAFE is almost perfectly aligned with human judgment than other evaluators (see Table 1). Similarly, individual LLM judges show more substantial to nearly perfect agreement with human judgments than EM and BERTScore.
Due to the high-class imbalance in TriviaQA, kappa scores can be misleadingly low despite high raw agreement - a known limitation called the “kappa paradox” (Cicchetti and Feinstein, 1990). Therefore, we treat the evaluation as a binary classification task where we consider each evaluator’s predictions against the human majority and report Macro-F1 scores which give equal weight to both classes regardless of their frequency in the selected random samples.
Table 1: Cohen’s Kappa scores displaying the agreement levels of various evaluators with human judgments across candidate models and tasks. Higher scores indicate better agreement with human judgments.
Evaluators
LLMs Tasks EM BS DeepSeek Llama GPT Mixtral Mistral DAFE
Llama AmbigQA 0.518 0.283 0.897 0.888 0.844 0.824 0.858 0.911
HotpotQA 0.577 0.498 0.885 0.877 0.899 0.820 0.832 0.953
NQ-Open 0.381 0.437 0.797 0.833 0.793 0.816 0.738 0.927
TriviaQA 0.281 0.564 0.460 0.547 0.439 0.396 0.299 0.684
GPT AmbigQA 0.561 0.252 0.951 0.944 0.897 0.861 0.853 0.967
HotpotQA 0.604 0.300 0.807 0.953 0.973 0.873 0.933 0.987
NQ-Open 0.453 0.218 0.809 0.884 0.824 0.824 0.829 0.956
TriviaQA 0.335 0.364 0.594 0.650 0.401 0.580 0.467 0.775
Mixtral AmbigQA 0.546 0.337 0.896 0.896 0.781 0.909 0.887 0.951
HotpotQA 0.546 0.349 0.920 0.940 0.933 0.859 0.940 0.973
NQ-Open 0.371 0.301 0.825 0.879 0.728 0.899 0.815 0.913
TriviaQA 0.317 0.390 0.661 0.625 0.605 0.678 0.436 0.764
Mistral AmbigQA 0.599 0.254 0.893 0.893 0.893 0.893 0.860 0.953
HotpotQA 0.605 0.383 0.903 0.937 0.902 0.895 0.937 0.958
NQ-Open 0.484 0.291 0.797 0.851 0.838 0.878 0.840 0.953
TriviaQA 0.467 0.239 0.754 0.758 0.725 0.645 0.470 0.854
As evidenced by consistently high Macro F1 scores in Table 2, DAFE maintains a strong alignment with human judgment. This represents a substantial improvement over individual model performance, where individual judges generally revealed varying levels of agreement with human evaluation. LLM-as-a-judge approach generally works better with larger more powerful models. This is particularly noticeable in DeepSeek and GPT which achieve higher Macro-F1 scores (0.97-0.98) across AmbigQA, HotpotQA, and NQ-Open compared to smaller models. This reveals an important scaling law in evaluation capability (Kaplan et al., 2020; Zheng et al., 2024; OpenAI et al., 2024). However, we also found that the most advanced models are not always guaranteed to be the best evaluators. We observed slightly comparable performance through small open-source Mistral-7B. For instance, when evaluating candidate Mixtral-8x7B on AmbigQA, Mistral-7B as-a-judge outperformed (0.944) judge GPT-3.5-turbo (0.891). Regardless, we observe relatively lower Macro-F1 scores for all LLM judges in TriviaQA.
Interestingly, despite EM’s deviation from the human majority (see Figure 2 and Table 6), lexical matching EM typically accomplishes better alignment with human evaluation on instance-level in Table 2 than neural-based BERTScore. EM’s strict and conservative nature leads to lower overall performance, but its high-precision characteristics ensure that when it identifies a match, it strongly aligns with human judgment. In contrast, BERTScore takes a more lenient approach to semantic matching. Although this leniency produces higher raw scores, it introduces more false positives, consequently reducing instance-level agreement with human judgments. This pattern emerges clearly in many models and tasks such as when evaluating Llama-3.1-70B on AmbigQA, EM shows a raw score of 42.3% but achieves a Macro-F1 of 0.744, while BERTScore indicates a higher raw score of 63.0% but a lower Macro-F1 of 0.641.
Table 2: Macro-F1 scores of various evaluators applied to different candidate LLMs and associated tasks. Higher scores indicate better performance. DAFE consistently achieves the highest Macro-F1 across all evaluated settings.
Evaluators
LLMs Tasks EM BS DeepSeek Llama GPT Mixtral Mistral DAFE
Llama AmbigQA 0.744 0.641 0.948 0.944 0.922 0.912 0.929 0.955
HotpotQA 0.778 0.745 0.942 0.939 0.949 0.910 0.916 0.976
NQ-Open 0.653 0.718 0.898 0.916 0.896 0.907 0.869 0.964
TriviaQA 0.612 0.782 0.726 0.772 0.717 0.695 0.640 0.842
GPT AmbigQA 0.792 0.622 0.976 0.972 0.949 0.930 0.927 0.984
HotpotQA 0.794 0.623 0.903 0.977 0.987 0.936 0.966 0.993
NQ-Open 0.703 0.606 0.904 0.942 0.911 0.911 0.914 0.978
TriviaQA 0.646 0.681 0.796 0.824 0.700 0.789 0.730 0.887
Mixtral AmbigQA 0.760 0.666 0.948 0.948 0.891 0.955 0.944 0.975
HotpotQA 0.761 0.657 0.960 0.970 0.966 0.930 0.970 0.987
NQ-Open 0.650 0.649 0.912 0.939 0.863 0.950 0.908 0.956
TriviaQA 0.625 0.695 0.829 0.812 0.803 0.838 0.716 0.882
Mistral AmbigQA 0.792 0.622 0.947 0.947 0.947 0.947 0.930 0.977
HotpotQA 0.796 0.673 0.951 0.969 0.951 0.947 0.969 0.979
NQ-Open 0.726 0.639 0.898 0.925 0.919 0.939 0.920 0.976
TriviaQA 0.718 0.608 0.925 0.879 0.863 0.822 0.735 0.927
Figure 3: Heatmap illustrating the performance of four individual LLM judges on HotpotQA. Each cell value represents percentages (%). Rows represent predicted outcomes (P: Positive, N: Negative), while columns represent actual outcomes. See Appendix C for full results.
4.2 Analysis
In our experiments, candidate LLMs generated 7,500 outputs for the given tasks, with each evaluator producing 7,500 corresponding evaluations. We randomly sampled 100 error cases (50 false positives and 50 false negatives) from each evaluator to understand their behavior. Given EM had only 11 false positives, we included all of them in our analysis. Due to space constraints, we moved the detailed analysis of EM and BERTScore to Appendix C and focused exclusively on the LLM-as-a-judge method here.
LLM-based evaluators demonstrate strong abilities in recognizing semantic variations while maintaining the core meaning, especially when assessing responses that use different terminology or structural approaches to convey the same information. For instance, in the evaluation examples, evaluators correctly identified that “Salma Hayek” and “Salma Hayek Pinault” refer to the same individual, acknowledging the semantic equivalence despite differences in phrasing. Similarly, when assessing responses that use different terms for the same entity, such as recognizing “Nick Fury, Agent of S.H.I.E.L.D.” as part of the broader “Marvel” universe, the evaluators effectively maintain the core meaning and contextual relevance. Their explanations show systematic assessment patterns that combine multiple evaluation criteria including factual accuracy, logical coherence, and contextual relevance.
LLMs are prone to hallucination in justification (Zhang et al., 2023), where they fabricate reasoning to support their evaluations, produce detailed but incorrect explanations, or reference non-existent criteria or standards. In LLM judges, false positives and negatives (e.g., see Figure 3) often result from overlooking critical distinctions between candidate LLM outputs and failing to account for the specificity required by the reference answer. This pattern is particularly noticeable in Mistral 7B, where the model disregards the ground truth and provides evaluations influenced by unknown factors. For example, when evaluating candidate GPT-3.5’s response “The foreign minister of Germany who signed the Treaty of Versailles was Hermann Müller.” which is correct according to the reference answer “Hermann Müller” and human evaluation, Mistral 7B as-a-judge incorrectly marked this response as false and fabricated reasoning “Hermann Müller was the Chancellor of Germany, not the Foreign Minister. The Foreign Minister of Germany who signed the Treaty of Versailles was Gustav Stresemann.” in support of its decision. The same problem can also be attributed to inconsistent evaluations. Because when Mistral 7B acted as a candidate for the same question, its response to the question is completely different: “The Treaty of Versailles was signed by Matthias Erzberger, a German politician who served as the President of the German National Assembly at the time”. There are also alternative interpretations of this issue, such as ambiguity in the question, but we leave a deeper exploration of these aspects to future work.
We observe a different pattern in some judges, specifically, GPT-3.5 and Mixtral 8x7B which focuses more on specificity. This approach shifts the evaluation towards false negatives by missing semantically similar but structurally different answers. We found many cases when such evaluators failed to account for valid variations in phrasing or granularity, focusing instead on rigid adherence to the reference answer. Compounding these issues are reasoning errors within the evaluators’ own explanations, which often contain fabrications, circular logic, or overconfident assertions. By insisting on correctness derived strictly from the reference, evaluators disregard valid alternative perspectives and can even mischaracterize or invert the facts in their attempts to justify their decisions. This dynamic leaves little room for nuance or ambiguity, and it pushes the evaluation process away from fair, context-sensitive assessment toward rigid, and sometimes inaccurate, verdicts.
Verbosity (Ye et al., 2024) emerges as a subtle source of bias, where more elaborate answers are sometimes overrated simply due to their detail and fluency, while concise yet correct responses are undervalued. This misplaced emphasis leads to irrelevant judgment criteria, such as praising the presence of irrelevant information or penalizing perfectly valid but succinct answers. We also found that LLM-based judges encounter challenges in multiple reference answers and more open-ended questions. This confusion is especially pronounced in the TriviaQA where the diversity and flexibility of valid responses present challenges for the judges’ ability to consistently recognize and evaluate a range of correct answers.
We found several temporal limitations in LLM-based evaluators. Although most of our datasets are older and the evaluator models are relatively up-to-date, we still observed instances where references to recent events, newly emerging terminology, or evolving contexts were misinterpreted. The FreshQA dataset (Vu et al., 2023), being recent, serves as a valuable testbed for assessing these temporal deficiencies. As shown in Table 3, LLM-based evaluators indicate deviation from human judgment on FreshQA compared to tasks that rely on older information, such as HotpotQA. Specifically, in dynamic or time-sensitive contexts, we found that LLM judges tend to hallucinate by consistently classifying candidate model responses as True, even when incorrect. For example, when presented with the question: “On what date did the Patriots last play the Miami Dolphins?” the LLM-generated response states: ‘‘The last time the New England Patriots played the Miami Dolphins was on January 1, 2023, during the NFL regular season.” Despite the correct reference answer being “November 24, 2024” the LLM evaluator not only failed to recognize the inaccuracy but also hallucinated an erroneous justification, stating: “The proposed answer correctly states the date the New England Patriots last played the Miami Dolphins as January 1, 2023, which matches the information provided.”
Table 3: Performance (in Macro F1) of LLM judges on FreshQA.
Evaluators
LLMs DeepSeek Llama GPT Mixtral Mistral DAFE
DeepSeek 0.714 0.692 0.715 0.614 0.724 0.830
Llama 0.801 0.835 0.737 0.817 0.730 0.917
GPT 0.659 0.695 0.824 0.780 0.746 0.891
Mixtral 0.732 0.708 0.779 0.738 0.703 0.936
Mistral 0.687 0.665 0.802 0.818 0.723 0.880
Figure 4: Disagreement rates between primary judges (Llama+Mistral) across candidate LLMs and tasks.
4.3 Disagreements between primary judges
Figure 4 shows that disagreements between our primary judges, Llama-3.1 70B and Mistral 7B, mainly occur in the TriviaQA and FreshQA, with disagreement rates reaching 20.3% and 44.3%, respectively. Interestingly, higher disagreement rates between primary judges create a greater opportunity for DAFE to refine evaluations. As depicted in Figure 4, FreshQA (31.3% for Llama-70B, 39.7% for Mistral-7B) demonstrates the highest disagreement, allowing DAFE to improve Macro F1 scores (see Table 3).
4.4 Impact of arbitration
Our proposed arbitration approach significantly enhanced evaluation performance by resolving disputes through an independent judge, GPT-3.5-turbo (see Figure 5). Notably, in the AmbigQA, Macro F1 scores advanced from 72.9% to 86.6%, and Cohen’s Kappa increased from 0.467 to 0.773 (see Figure 7). These improvements highlight the pivotal role of the arbitrator in ensuring reliable and consistent evaluation outcomes.
Figure 5: Comparison of Macro F1 scores before and after arbitration (see Appendix C for more results).
5 Related work
Evaluation of natural language generation has traditionally relied on supervised signal-based metrics such as EM which evaluates the exact lexical match between generated outputs and reference answers. Despite its simplicity and efficiency, EM overlooks semantically equivalent variations, often penalizing accurate responses that use different phrasing (Wang et al., 2024a; Kamalloo et al., 2023). Other commonly used metrics including BLEU Papineni et al. (2002) and ROUGE (Lin, 2004) primarily focus on n-gram overlap with human written reference texts. Despite their widespread use, these metrics have significant limitations in capturing semantic subtleties and contextual relevance (Zhang et al., 2020). To address the limitations of conventional metrics, various model-based methods such as BERTScore Zhang et al. (2020) offer semantically informed evaluation. However, even BERTScore and similar embedding-based methods struggle to effectively evaluate open-ended generation Zheng et al. (2024); Sun et al. (2022).
Recent advances in LLMs have unlocked new opportunities for automatic and context-aware evaluation (Li et al., 2024b; Chiang and Lee, 2023; Zheng et al., 2024). A key strength of LLM-based evaluators lies in their ability to operate in reference-free settings, where evaluation does not rely on pre-defined answers but instead leverages subjective criteria such as helpfulness, relevance, and coherence. This capability makes LLM evaluators particularly well-suited for assessing tasks where multiple valid responses exist or where human-like judgment is required (Li et al., 2024a). For instance, LLMs are frequently used in subjective evaluations such as pairwise comparison (“Which response is better?”) or single-response scoring (“How good is this response based on criteria X?” (Verga et al., 2024; Chan et al., 2024). LLM-based evaluators are specifically effective for tasks like summarization, where subjective criteria are central to evaluation (Liu et al., 2023). However, they are less effective for fact-based tasks such as free-form question-answering, where responses are either correct or incorrect and require explicit verification against reference answers.
Furthermore, LLM-based evaluators face several challenges, particularly in ensuring consistency and fairness (Ye et al., 2024; Khan et al., 2024). In reference-free settings, the absence of a definitive ground truth increases the risk of bias in evaluations (Ye et al., 2024; Kim et al., 2024; Huang et al., 2024a). Common biases include positional bias, where LLMs may favor responses based on their order (Zheng et al., 2024; Khan et al., 2024), verbosity bias, which favors longer or more detailed responses (Huang et al., 2024b), and self-enhancement bias, where models may disproportionately prefer their own outputs (Zheng et al., 2024). These biases can distort evaluations and undermine the reliability of the results.
6 Conclusion
We present DAFE, a framework designed to evaluate free-form question-answering by leveraging LLMs. Our findings demonstrate that individual LLM judges are reliable alternatives to traditional lexical and neural-based metrics, offering closer alignment with human evaluations. However, relying solely on individual judges poses challenges including inherent biases and prompt sensitivity, which can affect evaluation performance. DAFE addresses these challenges through a dynamic arbitration mechanism. This design achieves near-perfect agreement with human evaluations, establishing DAFE as a trustworthy and reliable framework for evaluating open-ended language generation tasks. In the future, we aim to explore DAFE by excluding reference answers and integrating LLM agents with tools-interacting capabilities for evaluation.
7 Limitations
We acknowledge certain limitations in our study. The accuracy of evaluations depends on the quality and clarity of reference answers, which serve as the basis for determining correctness. Inconsistent or ambiguous references could affect evaluation outcomes. Similarly, this study primarily uses binary verdicts which might overlook detailed aspects of responses that could be captured through more comprehensive evaluation criteria. Furthermore, while we conducted an error analysis of LLM judges and automatic metrics, there may be error cases that were not identified during our manual review, leaving gaps in understanding the full spectrum of evaluation inaccuracies. Finally, our study focuses exclusively on English, and the applicability of our approach to other languages, particularly morphologically rich or resource-scarce ones, remains unexplored.
Appendix A Free-form Question-Answering
In our experiments, we include AmbigQA (Min et al., 2020), FreshQA (Vu et al., 2023), HotpotQA (Yang et al., 2018), Natural Questions (Kwiatkowski et al., 2019), and TriviaQA (Joshi et al., 2017).
  • AmbigQA: Focuses on 14K ambiguous questions derived from NQ, requiring systems to identify multiple valid interpretations and generate disambiguated questions alongside corresponding answers.
  • FreshQA: A QA benchmark containing 600 questions that consist of a diverse range of types, including those requiring fast-changing world knowledge and questions with false premises that need debunking. It is regularly updated to reflect current information and is designed to evaluate the factual accuracy of LLMs in handling up-to-date and evolving knowledge.
  • HotpotQA: Contains 113K questions based on Wikipedia. It is designed to test multi-hop reasoning, requiring connections across multiple paragraphs, and includes annotated supporting facts for evaluation.
  • Natural Questions (NQ): Consists of real user queries from Google Search, paired with Wikipedia articles. The dataset includes 307K training examples annotated with both long (paragraph) and short (entity-level) answers.
  • TriviaQA: Features approximately 650K trivia questions, with evidence sourced from Wikipedia and web searches. These questions often require reasoning across multiple documents for complex answer synthesis.
We utilize the validation splits across multiple datasets: the standard validation split for AmbigQA and Natural Questions, the “distractor” subset’s validation split for HotpotQA, and the “unfiltered.nocontext” subset’s validation split for TriviaQA. We randomly sampled 300 examples from each dataset using Seed 42.
Appendix B Human evaluation
This section provides detailed guidelines for human annotators responsible for evaluating the outputs of candidate LLMs. The goal is to ensure consistency and objectivity across all evaluations. These guidelines provide clear instructions for assessing each model’s response based on its alignment with the reference answer and contextual relevance.
B.1 Guidelines
Dear Evaluator,
Thank you for your valuable contribution to this evaluation process. These guidelines outline the process for evaluating Large Language Model (LLM) outputs for the given tasks. As annotators, you will receive three components for each evaluation instance: the input question, reference answer(s), and the model’s response. Your task is to evaluate the responses independently and score them on a binary scale: ‘1’ for ‘True’ (correct) and ‘0’ for ‘False’ (incorrect).
A response warrants a score of ‘1’ when it demonstrates semantic equivalence with the reference answer, even if expressed through alternative phrasing or structure. This includes acceptable variations such as synonym usage and structural variations. Additional contextual information is acceptable as long as it doesn’t introduce errors.
Responses receive a score of ‘0’ when they contain factual errors, miss crucial elements from the reference answer, or demonstrate contextual misalignment. Partial answers that omit essential information should be marked incorrect, regardless of the accuracy of included content. When multiple reference answers are provided, a response is correct if it fully aligns with at least one reference.
You are encouraged to use internet resources when needed to verify specific facts, terminology, or potential synonyms that may affect your evaluation decision. However, the reference answer should remain the primary basis for evaluation. Focus on whether the model’s response conveys the same core information as the reference answer. To maintain reliability, document any challenging cases requiring further discussion with other annotators.
B.2 Inter human annotator agreement
We calculate Fleiss’ Kappa ($\kappa$(Fleiss and Cohen, 1973) to assess inter-rater reliability among human annotators. Table 4 and 5 show the inter-annotator agreement across models and tasks.
LLMs AmbigQA FreshQA HotpotQA NQ-Open TriviaQA
DeepSeek 0.975 0.949 0.986 0.889 0.456 ($\kappa$ paradox)
Llama 0.945 0.962 0.973 0.985 0.935
GPT 0.989 0.973 0.982 0.990 0.948
Mixtral 0.981 0.945 0.996 0.977 0.936
Mistral 0.978 0.932 0.981 0.978 0.975
Table 4: Fleiss’ Kappa scores of human annotators across models and tasks.
The results demonstrate high reliability, with Fleiss’ Kappa scores consistently above 0.93 for most tasks. The highest agreement is observed in Mixtral evaluations on HotpotQA ($\kappa=0.996$), and GPT on NQ-Open ($\kappa=0.990$). In FreshQA, which shows lower Kappa scores, the agreement among annotators remains high including 99.3% in GPT and 98.0% in Mixtral.
The percent agreement scores in Table 5 further confirm strong inter-annotator consistency. Most models achieve over 98% agreement across AmbigQA, HotpotQA, NQ-Open, and TriviaQA. However, DeepSeek exhibits lower agreement on NQ-Open (92.0%) and TriviaQA (90.0%). This indicates a variance in human ratings for these tasks.
LLMs AmbigQA FreshQA HotpotQA NQ-Open TriviaQA
DeepSeek 99.0% 98.0% 99.7% 92.0% 90.0%
Llama 96.3% 98.0% 98.0% 99.0% 99.0%
GPT 99.3% 99.3% 98.7% 99.3% 99.0%
Mixtral 98.7% 98.0% 99.7% 98.3% 98.3%
Mistral 98.3% 97.0% 98.7% 98.3% 99.0%
Table 5: Human annotators percent agreement scores across candidate models and tasks.
Appendix C Additional results
This section provides further results and analysis of conventional metrics and LLM-based evaluators. Table 6 illustrates the overall performance of candidate LLMs obtained through various evaluators. Unlike lexical matching and neural-based metrics, each LLM-as-a-judge indicates overall performance close to the human majority. Automatic metrics like EM severely underestimate the candidate LLMs’ performance. On the other hand, BERTScore tends to overestimate the performance.
EM underestimates performance because it requires a candidate’s response to exactly match one of the reference answers. This rigid, lexical approach fails to account for valid paraphrases, synonyms, or alternative expressions that convey the same meaning. In free-form QA tasks, where there can be multiple correct answers phrased in various ways, EM’s strict criteria often penalize responses that are semantically accurate but differ slightly in wording. As a result, it underestimates the true capabilities of candidate LLMs, leading to an incomplete assessment of their performance.
BERTScore relies on token-level semantic similarity, which rewards shallow lexical overlap rather than actual factual accuracy. For example, in cases where minor differences in wording (e.g., “The Treaty of Versailles was signed in 1919.” versus “The Treaty of Versailles ended in 1919.”) lead to opposing factual claims, BERTScore still scores the response high due to its emphasis on matching tokens (e.g., “signed” versus “ended”). Additionally, verbosity bias and threshold instability—where a default threshold (threshold = 0.5) is arbitrarily set—further inflate its raw accuracy. However, when comparing raw accuracy with instance-level agreement metrics like Cohen’s kappa, which adjusts for class imbalance and penalizes asymmetric errors, the limitations of BERTScore become apparent.
LLMs Tasks Evaluators
EM BS HM DeepSeek Llama GPT Mixtral Mistral
DeepSeek AmbigQA 56.3 80.0 84.3 86.3 73.7 75.0 62.3 93.3
FreshQA 31.3 88.0 84.3 84.7 82.7 75.3 58.0 82.3
HotpotQA 38.6 78.4 57.7 58.0 51.0 51.0 52.7 57.7
NQ-Open 35.0 78.3 60.3 64.7 63.7 61.3 55.3 68.3
TriviaQA 77.3 90.7 94.3 90.7 94.0 91.7 81.7 89.7
Llama AmbigQA 42.3 63.0 67.0 64.0 65.3 64.7 63.0 66.0
FreshQA 25.6 81.3 77.7 81.3 78.3 72.7 71.0 62.3
HotpotQA 34.3 67.7 56.3 56.7 58.3 54.0 50.7 52.7
NQ-Open 31.7 61.7 66.3 62.3 62.7 60.0 59.0 66.7
TriviaQA 74.3 94.0 94.7 88.0 90.3 90.0 88.7 84.7
GPT AmbigQA 49.7 78.0 71.7 70.3 70.0 68.0 65.7 71.0
FreshQA 24.6 89.3 70.7 58.0 51.7 78.7 83.0 83.3
HotpotQA 33.7 80.0 54.0 50.3 53.0 52.7 51.7 54.0
NQ-Open 36.3 74.0 65.3 65.3 62.7 59.0 59.0 67.0
TriviaQA 74.3 95.3 93.0 90.0 89.3 90.7 89.7 86.3
Mixtral AmbigQA 37.7 70.3 61.7 58.7 57.3 62.0 59.3 61.7
FreshQA 18.6 89.7 86.0 72.3 67.0 87.0 85.0 77.7
HotpotQA 25.0 69.7 47.0 46.3 45.3 45.7 44.7 46.0
NQ-Open 23.7 63.7 56.7 54.0 52.7 47.7 52.3 59.7
TriviaQA 64.7 91.3 90.7 83.7 86.3 89.7 86.0 85.3
Mistral AmbigQA 31.0 61.7 49.7 47.7 46.3 47.7 46.3 53.3
FreshQA 15.6 80.0 81.7 60.7 59.0 83.7 84.0 86.0
HotpotQA 23.7 64.7 40.0 39.3 39.0 38.0 37.0 39.0
NQ-Open 22.7 60.0 46.0 41.3 40.0 43.3 41.3 50.0
TriviaQA 62.0 94.3 83.7 78.0 81.3 81.0 79.7 85.0
Table 6: Raw performance of candidate LLMs across free-form QA tasks evaluated through various methods. HM represents Human Majority and BS denotes BERTScore.
C.1 Impact of arbitration on dispute resolution
Figure 6 illustrates the impact of arbitration on resolving disagreements between primary judges. Arbitration, facilitated by GPT-3.5 as the tiebreaker, consistently improves performance across all tasks, particularly in FreshQA and TriviaQA, where Macro F1 increases by up to 21.5 points. In contrast, tasks like AmbigQA and HotpotQA, where primary judges initially exhibit stronger agreement, show smaller but still meaningful improvements. This highlights the critical role of arbitration in enhancing agreement and achieving closer alignment with ground truth, especially in cases of significant disagreement among primary judges.
Notably, evaluations of DeepSeek-v3 exhibit higher disagreement between Llama-3.1-70B and Mistral-7B, particularly in FreshQA (28.3%) and AmbigQA (25.7%). From our analysis, we did not find strong evidence explaining why DeepSeek-v3 leads to higher disagreement between the primary judges.
Figure 6: Impact of arbitration on disagreements between primary judges. Note that we used Llama-3.1-70B and Mistra 7B as primary judges. GPT-3.5-turbo is only utilized when disagreements are found. The models given in the figure are candidate LLMs which generate outputs for the given tasks and are then evaluated through DAFE.
We observed substantial enhancements in Cohen’s Kappa scores across several tasks. For instance, as illustrated in Figure 7, in the AmbigQA Cohen’s Kappa increased from 0.881 to 0.911 for Llama. Similarly, in the same task, Cohen’s Kappa from 0.467 to 0.773 for candidate DeepSeek. These improvements demonstrate that the arbitration mechanism effectively enhances the reliability and consistency of evaluations, particularly in complex and ambiguous tasks where primary judges are more likely to disagree.
Some Cohen’s Kappa scores remain relatively low, particularly in FreshQA and DeepSeek-evaluated outputs. This is partially explained by the Kappa Paradox, where high agreement on extreme cases (e.g., clear correct/incorrect responses) and unbalanced class distributions can artificially lower the Kappa scores. In such cases, even when evaluators mostly agree, Cohen’s Kappa can appear lower than expected. Despite this, the arbitration process effectively mitigates inconsistencies, especially in tasks involving evolving knowledge and nuanced interpretations, such as FreshQA.
Figure 7: Comparison of Cohen’s kappa scores before and after arbitration (GPT-3.5-turbo as arbitrator). The performance is illustrated across candidate LLMs and tasks.
C.2 Cost analysis
Human evaluation is the gold standard for assessing LLM-generated responses, but it is expensive and time-consuming. In our setup, we employed three human annotators who volunteered their efforts. However, if these annotators were compensated based on standard annotation rates, the cost of evaluating such outputs would be significantly higher. On the other hand, GPT-3.5-turbo, acting as an arbitrator in DAFE, incurs a cost that depends on the number of arbitration cases. In our evaluations, GPT-3.5-turbo was invoked 1,318 times, with an estimated total cost of \$0.59, which increases to \$5.40 if a 2048 max token setting is used (see Table 7). Since GPT-3.5 is only invoked when primary judges disagree, this selective arbitration substantially reduces overall evaluation expenses while maintaining high reliability in assessments. Rather than relying on a single model for evaluation, this multi-model arbitration approach enhances trust by mitigating biases and weaknesses inherent in any individual model.
By invoking the arbitrator only when disagreements occur (rather than evaluating all responses), DAFE reduces arbitration usage by 82–94% compared to a majority-voting system. This leads to:
  • Over 90% fewer third-judge inferences, drastically lowering computational demand.
  • Up to 95% cost savings by avoiding redundant model evaluations.
  • Better scalability, making it practical for large-scale deployments
Table 7: Cost-efficiency analysis of DAFE: Summary of disagreement rates and tiebreaker usage across candidate models and tasks
Candidate LLMs Tasks Samples Disagreement Rates (%) Tiebreaker Usage
DeepSeek AmbigQA 300 25.7 77
FreshQA 300 28.3 85
HotpotQA 300 10.7 32
NQ-Open 300 12.0 36
TriviaQA 300 14.3 43
Llama AmbigQA 300 10.0 30
FreshQA 300 31.3 94
HotpotQA 300 13.0 39
NQ-Open 300 18.0 54
TriviaQA 300 17.0 51
GPT AmbigQA 300 7.0 21
FreshQA 300 44.3 133
HotpotQA 300 5.7 17
NQ-Open 300 13.0 39
TriviaQA 300 15.7 47
Mixtral AmbigQA 300 9.0 27
FreshQA 300 37.3 112
HotpotQA 300 4.7 14
NQ-Open 300 13.0 39
TriviaQA 300 17.0 51
Mistral AmbigQA 300 11.7 35
FreshQA 300 39.7 119
HotpotQA 300 6.0 18
NQ-Open 300 14.7 44
TriviaQA 300 20.3 61
Total 7500 1318
C.3 DeepSeek as the arbitrator
To assess the impact of using DeepSeek as the arbitrator in DAFE, we conducted experiments by replacing GPT-3.5-turbo with DeepSeek. We evaluated this setup using different candidate models across multiple tasks. Specifically, we tested GPT-3.5 on TriviaQA, DeepSeek on NQ-Open, and Llama on FreshQA. The primary judges remained Llama and Mistral, and arbitration was invoked only in cases of disagreement. Our findings indicate that DeepSeek as the arbitrator achieves strong performance, with Macro-F1 scores of 91.23 on TriviaQA, 79.11 on NQ-Open, and 0.914 on FreshQA.
C.4 Evaluating with one strong LLM-as-a-judge
While a single state-of-the-art evaluator can achieve strong performance in many cases, the dual-LLM framework remains critical for ensuring robustness, particularly in high-stakes or ambiguous scenarios.
To explore the potential of a more powerful single LLM, we evaluated GPT-3.5-turbo on HotpotQA and TriviaQA using GPT-4o as a judge. With this configuration, GPT-4o as the evaluator achieved a Macro-F1 score of 0.946 on HotpotQA, demonstrating its exceptional capability. However, the same GPT-4o judge achieved only 0.784 on TriviaQA, which falls short of DAFE’s performance of 0.887. This shows that even the most advanced models show inconsistencies when evaluating free-form QA. This is particularly critical in precision-sensitive domains where minor errors can have outsized consequences.
In such settings, DAFE’s ensemble approach acts as a safeguard. When employing DAFE with GPT-3.5-turbo as the arbitrator, we achieved an even higher Macro-F1 of 0.984 on HotpotQA, surpassing the performance of a single GPT-4o. Interestingly, when we experimented with DeepSeek as the arbitrator in DAFE, performance remained strong at 0.963 Macro-F1, indicating that DAFE’s benefits are not solely tied to a specific arbitrator model.
C.5 Majority voting-based evaluation
We conducted additional experiments utilizing a traditional majority voting approach for evaluating candidate LLM performance. In this setup, we employed three LLM judges of equal weight: Llama, GPT-3.5, and Mistral to evaluate candidate models generated response. For every evaluation instance, each judge provided an independent binary verdict (True or False). The final decision is determined through a simple majority vote across these three verdicts.
As presented in Table 8, DAFE matches or closely approaches the Macro F1 and Cohen’s Kappa scores of the three-judge majority across almost all tasks and candidate LLMs. For example, on HotpotQA, evaluating candidate Llama with DAFE achieves a Macro F1 of 97.6% (compared to 97.6% for majority voting) and a Cohen’s Kappa of 0.95, while for GPT-3.5 on AmbigQA, DAFE reaches a Macro F1 of 98.4% (versus 98.3% for majority voting), indicating a negligible performance difference. Even in high-disagreement tasks like TriviaQA, where the primary judges (e.g., Mistral) disagree 20.3% of the time, DAFE retains strong alignment (with a Macro F1 of 92.7 compared to 93.5 for majority voting). Minor deviations, such as the one observed for candidate Mixtral on TriviaQA (DAFE’s Macro F1 = 0.88 vs. 0.95 for majority voting), reflect rare instances where both the primary judges and the arbitrator make errors, yet these outliers are substantially outweighed by the computational savings offered by selective arbitration.
Table 8: Comparison between Majority Voting (Llama+GPT-3.5+Mistral) and DAFE (GPT-3.5 as arbitrator). For each candidate LLM and task, the table reports Macro F1 and Cohen’s Kappa scores under Majority Voting, the disagreement rate (in %), and the corresponding scores using DAFE.
Candidate LLM Task Majority Voting Disagreement (%) DAFE
Macro F1 Kappa Macro F1 Kappa
Llama AmbigQA 95.5 0.91 10.0 95.5 0.91
HotpotQA 97.6 0.95 13.0 97.6 0.95
NQ-Open 96.3 0.93 18.0 96.4 0.92
TriviaQA 84.1 0.68 17.0 84.2 0.68
GPT AmbigQA 98.3 0.97 7.0 98.4 0.96
HotpotQA 99.3 0.99 5.7 99.3 0.98
NQ-Open 97.8 0.96 13.0 97.8 0.95
TriviaQA 90.5 0.81 15.7 88.7 0.77
Mixtral AmbigQA 98.9 0.98 9.0 97.5 0.95
HotpotQA 98.6 0.97 4.7 98.7 0.97
NQ-Open 98.3 0.97 13.0 95.6 0.91
TriviaQA 95.0 0.90 17.0 88.2 0.76
Mistral AmbigQA 97.6 0.95 11.7 97.7 0.95
HotpotQA 97.9 0.96 6.0 97.9 0.95
NQ-Open 97.6 0.95 14.7 97.6 0.95
TriviaQA 93.5 0.87 20.3 92.7 0.85
C.6 Impact of prompt variations
The effectiveness and consistency of LLM-based evaluation are significantly influenced by prompt design. Variations in prompt structure, reasoning order, explanation requirements, and task-specific examples can lead to notable differences in model verdicts. To analyze the robustness of the LLM judges in free-form QA, we conducted ablation studies on different prompt variations using Mistral as the candidate model and GPT as the judge.
C.6.1 Consistency in judgment across multiple trials
LLMs generate random text even at a temperature of 0. To assess whether this affects evaluation consistency, we repeated the same evaluation task five times for 100 Mistral-generated responses for HotpotQA.
  • Verdict stability: GPT produced identical True/False verdicts in 100% of cases. This suggest that its binary decision-making process remains stable even across multiple trials.
  • Explanation variability: While verdicts remained consistent, the rationales and explanations provided by GPT across trials, often cited different supporting facts for the same judgment.
C.6.2 Few-shot vs. zero-shot prompting
We investigated the impact of few-shot prompting where we included three task-specific examples in the prompt to guide the judge’s decision-making process. We found that adding few-shot examples resulted in a 2% increase in Macro-F1 scores. However, few-shot prompting introduced rigid decision patterns—the model sometimes over-applied reasoning from the examples rather than adapting flexibly to novel cases. For instance, multi-hop reasoning cases from HotpotQA, the judge model consistently followed the structure of the provided examples, even when the correct reasoning required a different approach.
C.6.3 Explanation requirement: Binary verdict vs. justification-based evaluation
To test whether requiring the model to generate explanations alongside verdicts improves judgment reliability, we compared two settings:
  • Binary verdict-only evaluation: The model was instructed to provide only a True/False response without any explanation.
  • Justification-based evaluation: The model was required to explain its reasoning before delivering the final verdict.
We found that:
  • Higher verdict volatility in verdict-only mode: When explanations were removed, 13% of verdicts changed between repeated evaluations of the same responses.
  • Reduced alignment with human judgment: Cohen’s Kappa agreement with human annotators dropped from 0.95 to 0.72, highlighting that rationale-based prompts lead to more stable and accurate decisions.
C.6.4 Reason-first vs. verdict-first prompting
In the verdict-first approach, the model is instructed to provide a True/False answer before justifying its decision, whereas in the reason-first approach, the model is asked to generate reasoning first and then conclude with a verdict. Experimental results showed no significant difference in accuracy or agreement scores between these two formats.
C.7 G-Eval: reference-free evaluation of free-form question-answering
Existing LLM-based evaluators such as G-Eval (Liu et al., 2023) are designed for reference-free, subjective tasks (e.g., summarization, dialogue), where evaluation criteria (e.g., coherence, fluency) are inherently ambiguous and scored on Likert scales. These frameworks prioritize qualitative judgments rather than binary factual correctness. In contrast, DAFE is explicitly tailored for reference-dependent, objective evaluation in free-form QA, where answers are either factually correct or incorrect based on alignment with explicit ground-truth references.
To validate this distinction, we tailored G-eval based method to investigate the capability of LLM-as-a-judge in reference-free settings. In this setting, we modify the evaluation prompt by excluding the reference answer $r$ and directly prompted the evaluator model as $P=\{x,\bar{y}\}$ along with instructions such as correctness.
The performance of LLM-as-a-judge drastically changes in reference-free settings. Without access to the ground truth references, we observe a stark decline in evaluation capability across all models (see Table 9 and 10 values in blue). This systematic deterioration spans all tasks and model combinations, though its severity varies by context. HotpotQA, with its demands for complex reasoning, exemplifies this challenge most clearly. The substantial gap between reference-based and reference-free evaluation underscores the crucial role of reference answers in reliable assessment.
Candidate LLMs Tasks Evaluators
EM BERTScore Human Majority Llama-3.1-70B GPT-3.5-turbo Mixtral-8x7B Mistral-7B
Llama-3.1-70B AmbigQA 42.3 63.0 67.0 65.3 [83.3] 64.7 [84.7] 63.0 [76.0] 66.0 [80.3]
HotpotQA 34.3 67.7 56.3 58.3 [81.0] 54.0 [81.0] 50.7 [67.3] 52.7 [69.3]
NQ-Open 31.7 61.7 66.3 62.7 [89.0] 60.0 [89.3] 59.0 [81.0] 66.7 [81.0]
TriviaQA 74.3 94.0 94.7 90.3 [90.3] 90.0 [90.3] 88.7 [89.0] 84.7 [84.0]
GPT-3.5 AmbigQA 49.7 78.0 71.7 70.0 [79.0] 68.0 [81.0] 65.7 [79.0] 71.0 [84.3]
HotpotQA 33.7 80.0 54.0 53.0 [85.3] 52.7 [85.7] 51.7 [82.3] 54.0 [86.3]
NQ-Open 36.3 74.0 65.3 62.7 [83.7] 59.0 [90.7] 59.0 [87.0] 67.0 [89.7]
TriviaQA 74.3 95.3 93.0 89.3 [89.0] 90.7 [88.7] 89.7 [90.3] 86.3 [84.3]
Mixtral-8x7B AmbigQA 37.7 70.3 61.7 57.3 [74.7] 62.0 [82.3] 59.3 [79.7] 61.7 [80.7]
HotpotQA 25.0 69.7 47.0 45.3 [80.0] 45.7 [84.7] 44.7 [72.0] 46.0 [78.0]
NQ-Open 23.7 63.7 56.7 52.7 [81.7] 47.7 [90.3] 52.3 [85.7] 59.7 [89.7]
TriviaQA 64.7 91.3 90.7 86.3 [85.7] 89.7 [89.0] 86.0 [86.7] 85.3 [86.0]
Mistral-7B AmbigQA 31.0 61.7 49.7 46.3 [61.0] 47.7 [78.7] 46.3 [74.7] 53.3 [85.0]
HotpotQA 23.7 64.7 40.0 39.0 [64.3] 38.0 [83.3] 37.0 [62.0] 39.0 [77.0]
NQ-Open 22.7 60.0 46.0 40.0 [72.3] 43.3 [85.7] 41.3 [78.0] 50.0 [92.3]
TriviaQA 62.0 94.3 83.7 81.3 [80.7] 81.0 [81.0] 79.7 [80.7] 85.0 [84.7]
Table 9: Overall performance (Accuracy) of candidate LLMs across free-form QA tasks. Values [in blue] represent LLM-as-a-judge in the reference-free mood.
Candidate LLMs Tasks Evaluators
EM BERTScore Llama-3.1-70B GPT-3.5-turbo Mixtral-8x7B Mistral-7B
Llama-3.1-70B AmbigQA 0.744 0.641 0.944 [0.629] 0.922 [0.604] 0.912 [0.669] 0.929 [0.631]
HotpotQA 0.778 0.745 0.939 [0.628] 0.949 [0.574] 0.910 [0.665] 0.916 [0.640]
NQ-Open 0.653 0.718 0.916 [0.606] 0.896 [0.560] 0.907 [0.639] 0.869 [0.622]
TriviaQA 0.612 0.782 0.772 [0.772] 0.717 [0.628] 0.695 [0.678] 0.640 [0.633]
GPT-3.5 AmbigQA 0.792 0.622 0.972 [0.686] 0.949 [0.603] 0.930 [0.596] 0.927 [0.553]
HotpotQA 0.794 0.623 0.977 [0.566] 0.987 [0.521] 0.936 [0.543] 0.966 [0.494]
NQ-Open 0.703 0.606 0.942 [0.671] 0.911 [0.544] 0.911 [0.601] 0.914 [0.536]
TriviaQA 0.646 0.681 0.824 [0.817] 0.700 [0.690] 0.789 [0.760] 0.730 [0.701]
Mixtral-8x7B AmbigQA 0.760 0.666 0.948 [0.704] 0.891 [0.636] 0.955 [0.654] 0.944 [0.622]
HotpotQA 0.761 0.657 0.970 [0.587] 0.966 [0.470] 0.930 [0.582] 0.970 [0.577]
NQ-Open 0.650 0.649 0.939 [0.652] 0.863 [0.517] 0.950 [0.590] 0.908 [0.529]
TriviaQA 0.625 0.695 0.812 [0.800] 0.803 [0.754] 0.838 [0.818] 0.716 [0.725]
Mistral-7B AmbigQA 0.792 0.622 0.947 [0.730] 0.947 [0.627] 0.947 [0.628] 0.930 [0.523]
HotpotQA 0.796 0.673 0.969 [0.649] 0.951 [0.478] 0.947 [0.680] 0.969 [0.578]
NQ-Open 0.726 0.639 0.925 [0.652] 0.919 [0.515] 0.939 [0.597] 0.920 [0.433]
TriviaQA 0.718 0.608 0.879 [0.881] 0.863 [0.840] 0.822 [0.846] 0.735 [0.744]
Table 10: Performance (Macro F1) of various evaluators across candidate LLMs and tasks. Values [in blue] represent LLM-as-a-judge in the reference-free mode.
C.8 DAFE in multi-reference answers
DAFE explicitly accommodates multiple gold reference answers by incorporating all available references into the judge LLM’s prompt during evaluation. For datasets like AmbigQA and TriviaQA, where questions often have multiple valid answers (e.g., synonyms, rephrased answers, or alternative factual representations), DAFE aggregates all reference answers into the judge’s input prompt (e.g., concatenating them as a comma-separated list).
This design ensures that the judge evaluates the candidate’s output against the full spectrum of acceptable answers, mirroring the human evaluation protocol, where annotators are instructed to mark a response as correct if it aligns with any reference answer. However, as presented in our paper, LLM-based judges encounter challenges with multiple reference answers. This confusion is particularly evident in TriviaQA, where multiple reference answers introduce difficulties for the judges to recognize and evaluate a range of correct responses.
C.9 Analysis of automatic metrics
Figures 8, 9, 10, and 11 illustrate the fundamental trade-offs in automatic metrics. In TriviaQA, where multiple normalized reference answers exist, EM achieves impressive true positives (61.7-74.3%) compared to HotpotQA (23.0-34.3%) which contains single reference answers. EM’s near-zero false positives across tasks (0-0.7%) stem from its strict string matching – it only flags matches when answers are identical to references. Our error analysis found three primary causes of such rare false positives including preprocessing errors, where character normalization removes crucial distinctions, and reference ambiguities, where incomplete or ambiguous references lead to incorrect matches. Additionally, a semantic mismatch occurs when the EM incorrectly labels a prediction as true by matching text without considering its context. For instance, despite their different contextual meanings, EM wrongly marks a match between a model prediction of “1944” (describing the start of a war) and a reference answer containing “1944” (representing the end of the war).
EM string-matching guarantees high precision and makes EM particularly effective when exact wording is crucial, such as mathematical problems. However, its rigid criteria also result in substantial false negatives (17.0-34.7%). These false negatives primarily occur when the candidate LLM generates semantically correct responses that differ from references in format or expression. Common cases include synonym usage and paraphrases, structural variations in phrasing (e.g., “School of Medicine at Harvard” vs. “Harvard Medical School”), granularity discrepancies where answers differ in levels of detail from references (e.g., answering “British writer” instead of “William Shakespeare”), and partial matches that contain valid information but don’t exactly mirror the reference.
Unlike EM, BERTScore offers advantages in capturing semantic similarities. In TriviaQA, it gains high true positive rates (81.3-92.0%) with relatively low false positives (2.0-13.0%). BERTScore’s performance varies significantly across tasks and is influenced by its sensitivity to the threshold setting. In HotpotQA, where answers require multi-hop reasoning, true positives reach 36.0-50.3%, with an increase in false positives (17.7-29.7%). A similar pattern appears in NQ-Open, with true positives of 43.3-53.0% and false positives of 10.7-21.0%. Its tendency toward false positives indicates that relying solely on embedding similarity often accepts answers that are contextually related but factually incorrect. The false positives emerge through semantic drift (where similar embeddings yield false matches), contextual misalignment (where word meanings shift based on context), and threshold instability (where similarity cutoffs fail to distinguish subtle semantic differences). Additionally, false positives emerge due to the verbose responses where additional content artificially increases similarity scores.
Figure 8: Confusion matrices comparing the performance of automatic metrics (EM, BERTScore) and individual LLM judges on AmbigQA.
Figure 9: Confusion matrices comparing the performance of automatic metrics (EM, BERTScore) and individual LLM judges on HotpotQA.
Figure 10: Confusion matrices comparing the performance of automatic metrics (EM, BERTScore) and individual LLM judges on NQ-Open.
Figure 11: Confusion matrices comparing the performance of automatic metrics (EM, BERTScore) and individual LLM judges on TriviaQA.
Appendix D Prompting
In our main experiment, we performed zero-shot prompting in the following two stages.
D.1 Prompting Candidate LLMs
We prompted candidate LLMs (see Figure 12) to record generations for each task. We set the same role and prompt structure for each candidate model to ensure the reproducibility of our results. Figure 13 shows the candidate GPT-3.5-turbo response at zero temperature for the input given in Figure 12.
You are a helpful assistant. Which comic book was also written by the writer of Crossed?
Figure 12: Prompting candidate GPT-3.5-turbo to elicit outputs for HotpotQA.
Crossed was written by Garth Ennis. The Boys is a comic book series also written by Garth Ennis. So the comic book that was also written by the writer of Crossed is The Boys.
Figure 13: Candidate GPT-3.5-turbo response to the input given in Figure 12.
D.2 Prompting LLM Judges
We prompted LLMs-as-judges to perform the evaluation (see Figure 14). In Figure 15, judge Llama-3.1-70B evaluating candidate GPT-3.5-turbo.
You are a helpful assistant acting as an impartial judge. You will be given a Question and a Proposed Answer. Your task is to judge whether the Proposed Answer is correct by comparing it to the Reference Answer. If the Proposed Answer is correct, choose ’True’, otherwise choose ’False’. Provide a brief explanation for your decision. Question: Which comic book was also written by the writer of Crossed? Provided Answer: Crossed was written by Garth Ennis. The Boys is a comic book series also written by Garth Ennis. So the comic book that was also written by the writer of Crossed is The Boys. So the answer is: The Boys. Reference Answer: the Vertigo series “Preacher” Evaluation: Provide your response in the following format: Decision: [True/False] Explanation: [Your brief explanation]
Figure 14: Prompting judge Llama-3.1-70B for evaluating candidate GPT-3.5-turbo.
Decision: False Explanation: The Proposed Answer incorrectly identifies "The Boys" as the comic book written by the writer of Crossed. While it is true that Garth Ennis wrote Crossed, the Proposed Answer fails to consider other works written by Garth Ennis. The Reference Answer, "Preacher", is indeed another comic book series written by Garth Ennis, specifically for the Vertigo imprint.
Figure 15: Llama-3.1-70B-Judge verdict on the candidate GPT-3.5-turbo output.

### v2

@@
-DAFE: LLM-Based Evaluation Through Dynamic Arbitration for Free-Form Question-Answering
+CLEV: LLM-Based Evaluation Through Lightweight Efficient Voting for Free-Form Question-Answering
 Abstract
-Evaluating Large Language Models (LLMs) free-form generated responses remains a challenge due to their diverse and open-ended nature.
-Traditional supervised signal-based automatic metrics fail to capture semantic equivalence or handle the variability of open-ended responses, while human evaluation, though reliable, is resource-intensive.
-Leveraging LLMs as evaluators offers a promising alternative due to their strong language understanding and instruction-following capabilities.
-Taking advantage of these capabilities, we propose the Dynamic Arbitration Framework for Evaluation (DAFE), which employs two primary LLM-as-judges and engages a third arbitrator only in cases of disagreements.
-This selective arbitration prioritizes evaluation reliability while reducing unnecessary computational demands compared to conventional majority voting.
-DAFE utilizes task-specific reference answers with dynamic arbitration to enhance judgment accuracy, resulting in significant improvements in evaluation metrics such as Macro F1 and Cohen’s Kappa.
-Through experiments, including a comprehensive human evaluation, we demonstrate DAFE’s ability to provide consistent, scalable, and resource-efficient assessments, establishing it as a robust framework for evaluating free-form model outputs.
+Evaluating free-form Question-Answering (QA) remains a challenge due to its diverse and open-ended nature.
+Traditional automatic metrics fail to capture semantic equivalence or accommodate the variability of open-ended responses.
+Leveraging Large Language Models (LLMs) as evaluators offers a promising alternative due to their strong language understanding and instruction-following capabilities.
+We propose the Consensus via Lightweight Efficient Voting (CLEV), which employs two primary LLMs as judges and engages a third judge only in cases of disagreement.
+This approach prioritizes evaluation reliability while reducing unnecessary computational demands.
+Through experiments, including human evaluation, we demonstrate CLEV’s ability to provide consistent, scalable, and resource-efficient assessments, establishing it as a robust framework for evaluating LLMs on free-form QA.
 1 Introduction
-The rapid advancements in Large Language Models (LLMs) have propelled the field of natural language processing forward, yet their evaluation remains a challenge (Laskar et al., 2024).
-In particular, free-form model responses are difficult to evaluate because their correctness depends on understanding the broader context and underlying meaning (Si et al., 2021).
-Many benchmarks, such as MMLU (Hendrycks et al., 2021), often simplify evaluation by focusing on structured formats (e.g., multiple-choice questions) (Chen et al., 2024).
-Although effective for certain tasks, such methods rely on log probabilities assigned to predefined options, where the model selects the most likely answer, limiting the range of capabilities that can be assessed (Thakur et al., 2024).
-This structured approach fails to accommodate the complexity of free-form responses, where multiple valid answers exist (Chang et al., 2024).
-The rigid, predefined options in such evaluations not only limit the scope of assessment but also overlook the diversity of potential correct responses in free-form tasks (Li et al., 2023; Zhang et al., 2024).
-Automatic metrics including lexical matching, n-gram, and neural-based have been widely adopted as scalable solutions for the evaluation of free-form model outputs.
-Lexical matching methods such as Exact Match (EM) evaluate model predictions by assessing strict lexical alignment between generated outputs and reference answers.
-However, EM fails to account for semantically equivalent variations in phrasing.
-For instance, despite their equivalence, EM treats “nuclear weapon” and “atomic bomb” as incorrect.
-Similarly, n-gram-based metrics (Papineni et al., 2002; Lin, 2004) primarily assess surface-level similarity and often fail to capture semantic equivalence, particularly when lexical or structural diversity conveys the same underlying meaning (Zhu et al., 2023; Chen et al., 2021; Zhang et al., 2020).
-Neural-based metrics like BERTScore (Zhang et al., 2020) address such limitations by leveraging contextual embeddings to evaluate semantic similarity.
-However, BERTScore depends on reference quality (Liu et al., 2024) and struggles with domain adaptation and length variations (Zhu et al., 2023).
-Furthermore, continuous score provider metrics are difficult to interpret (Xu et al., 2023).
-The limitations in automatic metrics become particularly evident when evaluating instruction-tuned chat models (Doostmohammadi et al., 2024), which tend to produce verbose and diverse responses (Saito et al., 2023; Wang et al., 2024b).
-Contrary to automatic metrics, human evaluation provides a more transparent assessment (Chiang and Lee, 2023).
-However, despite being the “gold standard”, human evaluation is not without its limitations.
-LLMs’ growing complexity and scale have made recruiting and coordinating multiple human raters increasingly resource-intensive and time-consuming (Mañas et al., 2024).
-Furthermore, the reliability of human evaluation is additionally challenged by variations in rater expertise and inherent subjectivity that affect reproducibility (Clark et al., 2021; Chiang and Lee, 2023).
-Recently, a paradigm shift has emerged where LLMs are utilized to judge the candidate model generations for given tasks (Zheng et al., 2024).
-This model-based method leverages the instruction-following capabilities of LLMs through evaluation prompts or, in some cases, fine-tuned versions of LLMs that are specifically optimized for evaluation.
-In this new line of work, research primarily focuses on pairwise comparison (Zheng et al., 2024; Wang et al., 2023; Vu et al., 2024), such as instructing an LLM to judge “which assistant response is better”, and single-answer scoring (Verga et al., 2024) like evaluating summarization task based on predefined criteria (e.g., likability, relevance, etc.) (Chiang and Lee, 2023; Hu et al., 2024; Liu et al., 2023; Chan et al., 2024; Chu et al., 2024).
-Inspired by a recent study on self-correction where external feedback helps models identify and correct their mistakes (Gou et al., 2024a), we propose to guide LLM-as-a-judge with human-annotated task-specific reference answers in order to explore the potential of LLMs as an alternative to lexical matching (e.g., EM), neural-based (e.g., BERTScore), and human evaluation for automatic evaluation of free-form model responses.
-Unlike traditional metrics, an LLM judge can leverage its language understanding and instruction-following capabilities to recognize the correctness of open-ended generations.
-We propose the Dynamic Arbitration Framework for Evaluation (DAFE), which employs LLM judges to evaluate free-form model responses.
-Using a single LLM as a judge, while simple, often leads to inconsistent evaluations, undermining trust in the results.
-On the other hand, the common practice of using large, universally capable models such as GPT-4 as evaluators makes the evaluation process both slow and costly (Jung et al., 2024; Adlakha et al., 2024; Verga et al., 2024), further limiting its broader applicability.
-Relying on multiple judges for every evaluation, though more reliable, exacerbates these computational challenges, making such approaches impractical at scale.
-DAFE offers a middle ground between these approaches by utilizing two complementary primary judges to perform the initial assessment.
-Only when these judges disagree, is a third independent arbitrator engaged to resolve the conflict.
-This selective arbitration ensures evaluation reliability and fairness while reducing computational overhead.
-Our experiments reveal that DAFE achieves significant improvements in metrics such as Macro F1 and Cohen’s kappa.
-Our key contributions include: a detailed analysis of limitations in conventional metrics for free-form QA, an evaluation of LLM judges with insights into their strengths and errors, a comprehensive human evaluation for benchmarking, and the introduction of DAFE—a scalable framework that improves reliability while minimizing the need for additional evaluators through selective arbitration.
-Figure 1: Our proposed Dynamic Arbitration Framework for Evaluation (DAFE).
-Two primary judges, $J_{1}$ and $J_{2}$, first provide verdicts $V_{i_{1}}$ and $V_{i_{2}}$ for an instance $i$.
-If agree, that consensus $V_{i}$ is the final decision $D_{i}$.
-If disagree, a tiebreaker model $J_{t}$ independently produces a verdict $V_{t}$.
-The final decision $D_{i}$ is then determined via majority voting among $\{V_{i_{1}},V_{i_{2}},V_{t}\}$.
+One of the practical limitations in evaluating free‑form Question‑Answering (QA) is the lexical–semantic mismatch between Large Language Model (LLM) outputs and the reference answers.
+For the query “Who wrote 1984?”, a dataset may list simply “George Orwell”, while a helpful model replies: “It was penned by the British author Eric Arthur Blair.” Although both refer to the same person, there is no surface token overlap, so lexical-matching and even embedding-based metrics assign it an unduly low score.
+For instance, Exact Match (EM) requires strict lexical alignment (e.g., failing to equate “nuclear weapon” and “atomic bomb”) and ignores semantic equivalence (doostmohammadi2024reliable).
+A reference-aware LLM-as-a-judge can instead reason over meaning, recognize that the candidate entails the gold fact, and deliver a verdict, thereby overcoming this lexical–semantic gap (10.5555/3666122.3668142; verga2024replacing).
+Existing studies using LLM as judges primarily focus on subjective pairwise comparison (wang2023large; vu2024autoraterstaminglarge) and single-answer scoring (chianglee2023large; hu2024llmbased; liuetal2023g).
+However, objective evaluation using LLM judges, particularly for free-form QA, has received limited attention.
+Furthermore, LLM-based judging also lies on a cost–quality spectrum.
+Querying a single judge is efficient but is less reliable due to the known limitations, such as prompt sensitivity, inconsistency, and bias (ye2024justiceprejudicequantifyingbiases), as no individual model captures the full diversity of reasoning styles, long-tail knowledge, and user values (feng2025llmdroolsmultillmcollaboration).
+To improve robustness, some studies employ fixed ensembles of three or more LLM judges and aggregate their decisions via majority voting (badshah-sajjad-2025-reference; verga2024replacing).
+While this improves reliability, it increases computational cost and latency (2024quantifyingcapabilitiesllmsscale), limiting scalability for large-scale evaluation (jung2024trustescalatellmjudges; adlakha2024evaluating).
+To address these trade-offs, we propose the Consensus via Lightweight Efficient Voting (CLEV) that balances the reliability and efficiency of using LLMs as judges for free-form QA.
+CLEV employs two primary judges for initial assessments and invokes a third judge only when disagreements occur.
+By minimizing redundant calls in the fixed three-judge majority vote setting, CLEV reduces computational overhead by roughly 80 to 95% (varying by task) while achieving substantial to perfect agreement.
+Our key contributions include: 1) establishing a principled LLM-as-a-judge evaluation setup for free-form QA that moves beyond token-level string matching metrics, 2) proposing CLEV, a lightweight consensus-based evaluation method that preserves the reliability of multi-judge voting while significantly reducing computational cost, 3) empirically validating CLEV across diverse QA datasets and multiple state-of-the-art LLMs, and 4) systematic analysis of failure cases.
+Figure 1: Our proposed Consensus via Lightweight Efficient Voting.
 2 Methodology
-This section briefly describes the key components of our proposed framework.
-2.1 Candidate LLMs
-A candidate LLM $\mathcal{C}_{\text{llm}}$ generates output $\bar{y}$ for the given input $x$. We first utilized candidate LLMs to obtain outputs for the given free-form question-answering tasks.
-2.2 LLMs-as-a-Judge
-A judge $\mathcal{J}_{\text{llm}}$ LLM delivers evaluation or verdict $V$ on candidate LLMs $\mathcal{C}_{\text{llm}}$ outputs $\bar{y}$.
-The $\mathcal{J}_{\text{llm}}$ evaluates output when prompted with $x$ (i.e., $x\rightarrow\mathcal{A}_{\text{llm}}$) and $\bar{y}$.
-We utilized the reference answer $r$ and prompted $P$ the $\mathcal{J}_{\text{llm}}$ as:
-$P=\{x,\bar{y},r\}$
-Utilizing $P$, $\mathcal{J}_{\text{llm}}$ performs the evaluation and delivers a decision as $V=J(P)$.
-The structure of this $V$ depends on the instructions provided in $P$.
-For instance, if a binary $V$ is required, $J$ assesses whether $\bar{y}$ is aligned with $r$ given the context $x$ and returns True if $\bar{y}$ is deemed correct, or False if it is not.
-The evaluation $P$ may vary from zero-shot, where $\mathcal{J}_{\text{llm}}$ receives no prior examples, to few-shot, which includes several related examples, or a chain of thought, encouraging $\mathcal{J}_{\text{llm}}$ to reason stepwise through the problem.
-2.3 Dynamic Arbitration Framework for Evaluation (DAFE)
+This section outlines the core components of our evaluation framework.
+Given a question and a candidate model’s answer, the evaluation task is to determine whether the answer is factually correct with respect to a reference.
+Because free-form QA requires an objective correctness decision rather than scalar scoring, we prompt LLM judges to return a binary verdict along with a brief rationale explaining their judgment.
+The verdict serves as the evaluation signal, while the rationale provides interpretability and supports later analysis.
+2.1 Consensus via Lightweight Efficient Voting (CLEV)
 In traditional human evaluation settings, when two annotators disagree on a judgment, a third expert is often called upon to resolve the dispute.
-Drawing inspiration from this efficient human arbitration practice, we propose the Dynamic Arbitration Framework for Evaluation (DAFE).
-Rather than immediately employing a large powerful or a closed-source LLMs-as-a-judge, DAFE adopts a cost-efficient approach by beginning with two complementary open-source models as primary judges based on their past performance (Kenton et al., 2024).
+Drawing inspiration from this efficient practice, we propose CLEV.
+As illustrated in Figure 1, rather than immediately employing three LLMs, CLEV adopts an efficient approach by beginning with two models as primary judges.
 When these judges reach a consensus, no further evaluation is needed.
-Only in cases of disagreement is the more powerful LLM engaged as an arbitrator, whose decision then creates a majority verdict.
-This dynamic approach maintains evaluation quality while minimizing reliance on expensive models.
-The method also accounts for varying skill levels across different LLMs and tasks (Liang et al., 2024; Sun et al., 2024).
-Formally, let $V_{i_{1}}$ and $V_{i_{2}}$ denote the verdicts from the two primary judges for the $i$-th evaluation instance. We define the agreement status $A_{i}$ as:
-$A_{i}=\begin{cases}1&\text{if }V_{i_{1}}=V_{i_{2}},\\[6.0pt] 0&\text{otherwise}.\end{cases}$
-If $A_{i}=1$, the final decision $D_{i}$ is simply $V_{i}$, the agreed-upon verdict of the primary judges.
-If $A_{i}=0$, a tiebreaker model provides an additional verdict $V_{t}$.
-The final decision $D_{i}$ is then obtained via majority voting among $\{V_{i_{1}},V_{i_{2}},V_{t}\}$.
-Formally:
-$D_{i}=\begin{cases}V_{i}&\text{if }A_{i}=1,\\[6.0pt] \text{majority}(\{V_{i_{1}},V_{i_{2}},V_{t}\})&\text{if }A_{i}=0.\end{cases}$
-The majority operation selects the verdict that appears at least twice among $\{V_{i_{1}},V_{i_{2}},V_{t}\}$. Since there are three votes, at least two must coincide for a majority.
+Only in cases of disagreement, the third LLM is engaged, whose decision then creates a majority verdict.
+2.2 Judges inclusion and exclusion criteria
+To systematically select suitable judges for CLEV, we evaluate various LLMs (see Figure 2) using 100 random instances from HotpotQA.
+For each model, we compare binary verdicts against human annotations and compute Cohen’s Kappa ($\kappa$) and Macro F1.
+We interpret $\kappa$ following the commonly used guideline where values between 0.610.610.800.80 indicate substantial agreement, and values above $0.80$ indicate near-perfect agreement (mchugh2012interrater).
+However, since $\kappa$ is known to be sensitive to class imbalance (cicchetti1990high), we together consider Macro F1 to ensure balanced evaluation across both classes:
+$\text{status}=\begin{cases}\text{primary judges},&\kappa\geq 0.6\ \text{and}\ \text{F1}\geq 0.85,\\[4.0pt] \text{third judge},&\kappa\geq 0.8\ \text{and}\ \text{F1}\geq 0.9,\\[4.0pt] \text{excluded},&\text{otherwise}.\end{cases}$
+Figure 2: Judges selection based on the defined criteria in Section $2.$
 3 Experiments
-We utilize the following settings to examine the performance and reliability of individual LLM judges and DAFE.
-3.1 Models
-We select open and closed-source instruct models to serve as candidates and judges in our experiment.
-These include DeepSeek-V3 (DeepSeek-AI et al., 2025), Llama-3.1 70B [^1] https://huggingface.co/meta-llama/Meta-Llama-3.1-70B-Instruct (Meta AI, 2024), GPT-3.5-turbo (Brown et al., 2020), Mistral 7B [^2] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3 (Jiang et al., 2023), and Mixtral 8x7B [^3] https://huggingface.co/mistralai/Mixtral-8x7B-Instructv0.1 (Jiang et al., 2024).
-We also utilize GPT-4o (OpenAI et al., 2023) and DeepSeek-R1 (DeepSeek-AI et al., 2025) in our ablation experiments.
-To ensure the reproducibility of our experiments, we set the temperature to 0 for all models under study, as the performance of LLM-based evaluators has been shown to drop when temperature increases (Hada et al., 2024).
-For our proposed DAFE method, we utilized Mistral 7B and Llama 3.1 70B as primary judges with GPT-3.5-turbo as the tiebreaker.
-In addition, we experiment with other models as tiebreakers in our ablation experiments.
-In the rest of the paper, we refer both candidate and judge LLMs as: DeepSeek, Llama, GPT, Mistral, and Mixtral.
-3.2 Datasets
-We focus on free-form question-answering (QA) since it has widespread practical applications and the critical importance of truthfulness in this domain (Gou et al., 2024a; Evans et al., 2021).
-In our experiment, we utilize five free-form QA datasets: AmbigQA (Min et al., 2020), FreshQA (Vu et al., 2023), HotpotQA (Yang et al., 2018), Natural Questions (Kwiatkowski et al., 2019), and TriviaQA (Joshi et al., 2017).
+We utilize the following settings to evaluate CLEV:
+Models
+Our candidate models includes Llama-3.1 70B (MetaLlama3_2024), GPT-3.5-turbo (NEURIPS2020_1457c0d6), Mistral 7B (jiang2023mistral), and Mixtral 8x7B (jiang2024mixtralexperts).
+For judges, Mistral‑7B met our inclusion criteria ($\kappa\geq 0.6$, $\mathrm{F1}\geq 0.85$) while offering low cost and fast inference.
+Llama‑3.1‑70B, though more expensive, provided substantial agreement.
+We therefore use both as primary judges to capture a range of capabilities (feng2025llmdroolsmultillmcollaboration; liangetal2024encouraging; sun2024skillaggregation).
+GPT‑3.5‑turbo serves as the third judge because it exceeds the required thresholds ($\kappa\geq 0.8$, $\mathrm{F1}\geq 0.9$).
+All models are run with temperature 0 to ensure reproducibility, as higher temperatures degrade evaluator reliability (hada2024llmbasedevaluators).
+Datasets
+We utilize four datasets: AmbigQA (minetal2020ambigqa), HotpotQA (yang2018hotpotqa), Natural Questions (kwiatkowskietal2019), TriviaQA (joshi2017triviaqa).
+We only utilize FreshQA to evaluate the judge’s ability to detect outdated information.
 See Appendix A for details.
-3.3 Prompts
-We designed generalized (i.e., with minimum instructions) zero-shot prompts with role-playing (Kong et al., 2024) for both candidates and judges.
-Initially, we prompt candidate LLMs to elicit outputs for the given random samples associated with each dataset.
-To evaluate the outputs of candidate LLMs, we prompt judge LLMs for binary verdicts (i.e., True or False) using $P=\{x,\bar{y},r\}$ and instructed to provide a brief explanation for their verdicts (see Appendix D for examples).
-Binary verdicts explicitly differentiate between correct and incorrect answers, minimize subjective interpretations, and simplify the evaluation process, thus facilitating automatic evaluation.
-In addition to three key prompt components (i.e., $x$, $\bar{y}$, $r$), we define the role of the judge LLMs as “You are a helpful assistant acting as an impartial judge.” to mitigate biases in judgments (Zheng et al., 2024).
-We chose not to use few-shot or chain-of-thought prompting strategies to keep the solution robust to a variety of tasks.
-Previous studies have also shown that in-context examples do not significantly improve the performance of model-based evaluators (Hada et al., 2024; Min et al., 2022).
-3.4 Baselines
-We establish the following baselines.
-Exact Match (EM):
-For our selected datasets and also free-form QA tasks, EM serves as a standard lexical matching metric to evaluate candidate LLM performance (Izacard and Grave, 2021; Lewis et al., 2020; Gou et al., 2024b).
-Due to the verbose nature of LLM-generated responses, we adapt EM to classify an answer as correct if any golden answer $r_{i}\in R$ appears within the generated response $\bar{y}$ (i.e., $r_{i}\subseteq\bar{y}$), rather than requiring complete strict string equality (i.e., $\bar{y}=r_{i}$).
-BERTScore:
-We use BERTScore (Zhang et al., 2020) which measures similarity by comparing contextualized word embeddings derived from a pre-trained BERT model.
-This enables the evaluation to focus on semantic correctness rather than exact lexical matches.
-As BERTScore is based on continuous values between -1 and 1, we set a threshold of $\tau=0.5$ to convert continuous similarity scores into binary 0 and 1.
-The purpose of this conversion is to allow direct comparison with other evaluation methods.
-For our implementation, we use the microsoft/deberta-xlarge-mnli [^4] https://huggingface.co/microsoft/deberta-xlarge-mnli model (He et al., 2021).
-G-Eval:
-In addition to automatic metrics, we also utilize G-Eval (Liu et al., 2023), a reference-free framework that uses GPT-4 to assess the quality of the generated text.
-In this setting, we modify the evaluation prompt by excluding the reference answer $r$ and directly prompted the evaluator model as $P=\{x,\bar{y}\}$ along with instructions.
-Human Evaluation:
-It remains the gold standard for assessing the outputs of candidate LLMs.
-We recruit three graduate students from our academic network, all specialized in natural language processing, to serve as annotators.
-We provide the input given to the candidate LLMs, reference answers, and candidate LLMs responses.
-This format, while similar, is distinct from the judge LLMs prompts which additionally require formatted decisions.
-We anonymize the origin of model responses to reduce potential bias linked to model familiarity or reputation.
-The annotators were asked to score the candidate LLMs outputs on a binary scale: ‘1’ for ‘True’ and ‘0’ for ‘False’ based on alignment with the reference answer and contextual relevance.
-For inter-rater reliability, we compute Fleiss’ Kappa ($\kappa$(Fleiss and Cohen, 1973) and percent agreement.
-See Appendix B for details.
+Prompts
+We design minimal zero-shot role-playing prompts (kong2024groleplay) for both candidate and judge LLMs.
+The judge LLM is instructed to return a binary verdict with a brief explanation, which reduces subjectivity and simplifies automatic evaluation (see Appendix F).
+Baselines
+We compare individual LLM judges and CLEV against EM, BERTScore (ZhangKWWA20) computed with microsoft/deberta-xlarge-mnli and thresholded at $\tau=0.5$ to yield binary decisions and majority voting, which always invokes a fixed setup of LLM judges and returns the model verdict (see Appendix B).
+Human evaluation
+We invite three volunteer graduate students from our lab to act as annotators.
+We provided the questions, reference answers, and candidate LLM responses, without information about model identity, to avoid bias.
+Each response is scored on a binary scale based on correctness and relevance (see Appendix C).
+Evaluation metrics
+We compute Fleiss’ Kappa ($\kappa$(fleiss1973equivalence) and percent agreement to measure inter-rater reliability among human annotators.
+To compare evaluators with humans, we calculate Cohen’s Kappa (mchugh2012interrater) against the human majority on an instance level.
+Because of class imbalance, kappa can appear low despite high agreement (cicchetti1990high).
+Therefore, we also frame the task as binary classification and report Macro-F1 scores.
 4 Results
-Figure 2 illustrates the raw performance of Llama obtained through various evaluators.
-Unlike lexical matching and neural-based metrics, each LLM-as-a-judge shows overall performance close to the human majority.
-The proposed DAFE method consistently achieves comparable or slightly better alignment with the human majority.
-Conventional metrics such as EM severely underestimate the candidate LLMs’ performance.
-Contrarily, BERTScore tends to overestimate the performance except in some cases such as when evaluating Llama on AmbigQA and NQ-Open (see Table 6 in Appendix C for additional results).
-Figure 2: Raw accuracy of candidate Llama across free-form QA tasks using automatic metrics and model-based evaluation. The Human Majority (HM) serves as the ground truth for all evaluators.
-4.1 Alignment with human evaluation
-We calculate Cohen’s kappa (McHugh, 2012) to find the agreement between each evaluator and the human majority to obtain instance-level comparison.
-Overall, DAFE is almost perfectly aligned with human judgment than other evaluators (see Table 1).
-Similarly, individual LLM judges show more substantial to nearly perfect agreement with human judgments than EM and BERTScore.
-Due to the high-class imbalance in TriviaQA, kappa scores can be misleadingly low despite high raw agreement - a known limitation called the “kappa paradox” (Cicchetti and Feinstein, 1990).
-Therefore, we treat the evaluation as a binary classification task where we consider each evaluator’s predictions against the human majority and report Macro-F1 scores which give equal weight to both classes regardless of their frequency in the selected random samples.
-Table 1: Cohen’s Kappa scores displaying the agreement levels of various evaluators with human judgments across candidate models and tasks.
-Higher scores indicate better agreement with human judgments.
-Evaluators
-LLMs Tasks EM BS DeepSeek Llama GPT Mixtral Mistral DAFE
-Llama AmbigQA 0.518 0.283 0.897 0.888 0.844 0.824 0.858 0.911
-HotpotQA 0.577 0.498 0.885 0.877 0.899 0.820 0.832 0.953
-NQ-Open 0.381 0.437 0.797 0.833 0.793 0.816 0.738 0.927
-TriviaQA 0.281 0.564 0.460 0.547 0.439 0.396 0.299 0.684
-GPT AmbigQA 0.561 0.252 0.951 0.944 0.897 0.861 0.853 0.967
-HotpotQA 0.604 0.300 0.807 0.953 0.973 0.873 0.933 0.987
-NQ-Open 0.453 0.218 0.809 0.884 0.824 0.824 0.829 0.956
-TriviaQA 0.335 0.364 0.594 0.650 0.401 0.580 0.467 0.775
-Mixtral AmbigQA 0.546 0.337 0.896 0.896 0.781 0.909 0.887 0.951
-HotpotQA 0.546 0.349 0.920 0.940 0.933 0.859 0.940 0.973
-NQ-Open 0.371 0.301 0.825 0.879 0.728 0.899 0.815 0.913
-TriviaQA 0.317 0.390 0.661 0.625 0.605 0.678 0.436 0.764
-Mistral AmbigQA 0.599 0.254 0.893 0.893 0.893 0.893 0.860 0.953
-HotpotQA 0.605 0.383 0.903 0.937 0.902 0.895 0.937 0.958
-NQ-Open 0.484 0.291 0.797 0.851 0.838 0.878 0.840 0.953
-TriviaQA 0.467 0.239 0.754 0.758 0.725 0.645 0.470 0.854
-As evidenced by consistently high Macro F1 scores in Table 2, DAFE maintains a strong alignment with human judgment.
-This represents a substantial improvement over individual model performance, where individual judges generally revealed varying levels of agreement with human evaluation.
-LLM-as-a-judge approach generally works better with larger more powerful models.
-This is particularly noticeable in DeepSeek and GPT which achieve higher Macro-F1 scores (0.97-0.98) across AmbigQA, HotpotQA, and NQ-Open compared to smaller models.
-This reveals an important scaling law in evaluation capability (Kaplan et al., 2020; Zheng et al., 2024; OpenAI et al., 2024).
-However, we also found that the most advanced models are not always guaranteed to be the best evaluators.
-We observed slightly comparable performance through small open-source Mistral-7B.
-For instance, when evaluating candidate Mixtral-8x7B on AmbigQA, Mistral-7B as-a-judge outperformed (0.944) judge GPT-3.5-turbo (0.891).
-Regardless, we observe relatively lower Macro-F1 scores for all LLM judges in TriviaQA.
-Interestingly, despite EM’s deviation from the human majority (see Figure 2 and Table 6), lexical matching EM typically accomplishes better alignment with human evaluation on instance-level in Table 2 than neural-based BERTScore.
-EM’s strict and conservative nature leads to lower overall performance, but its high-precision characteristics ensure that when it identifies a match, it strongly aligns with human judgment.
+In this section, we briefly report the results and refer the readers to Appendix E for detailed results.
+Evaluators (Cohen’s $\kappa$ [Macro-F1] vs Human Majority)
+Candid. Tasks EM BERTScore Llama GPT Mistral MV CLEV Disagr. (%)
+Llama AmbigQA 0.52 [0.74] 0.28 [0.64] 0.89 [0.94] 0.84 [0.92] 0.86 [0.93] 0.91 [0.96] 0.91 [0.96] 10.0
+HotpotQA 0.58 [0.78] 0.50 [0.75] 0.88 [0.94] 0.90 [0.95] 0.83 [0.92] 0.95 [0.98] 0.95 [0.98] 13.0
+NQ-Open 0.38 [0.65] 0.44 [0.72] 0.83 [0.92] 0.79 [0.90] 0.74 [0.87] 0.93 [0.96] 0.92 [0.96] 18.0
+TriviaQA 0.28 [0.61] 0.56 [0.78] 0.55 [0.77] 0.44 [0.72] 0.30 [0.64] 0.68 [0.84] 0.68 [0.84] 17.0
+GPT AmbigQA 0.56 [0.79] 0.25 [0.62] 0.94 [0.97] 0.90 [0.95] 0.85 [0.93] 0.97 [0.98] 0.96 [0.98] 7.0
+HotpotQA 0.60 [0.79] 0.30 [0.62] 0.95 [0.98] 0.97 [0.99] 0.93 [0.97] 0.99 [0.99] 0.98 [0.99] 5.7
+NQ-Open 0.45 [0.70] 0.22 [0.61] 0.88 [0.94] 0.82 [0.91] 0.83 [0.91] 0.96 [0.98] 0.95 [0.98] 13.0
+TriviaQA 0.34 [0.65] 0.36 [0.68] 0.65 [0.82] 0.40 [0.70] 0.47 [0.73] 0.81 [0.90] 0.77 [0.89] 15.7
+Mixtral AmbigQA 0.55 [0.76] 0.34 [0.67] 0.90 [0.95] 0.78 [0.89] 0.89 [0.94] 0.98 [0.99] 0.95 [0.98] 9.0
+HotpotQA 0.55 [0.76] 0.35 [0.66] 0.94 [0.97] 0.93 [0.97] 0.94 [0.97] 0.97 [0.99] 0.97 [0.99] 4.7
+NQ-Open 0.37 [0.65] 0.30 [0.65] 0.88 [0.94] 0.73 [0.86] 0.82 [0.91] 0.97 [0.98] 0.91 [0.96] 13.0
+TriviaQA 0.32 [0.63] 0.39 [0.70] 0.63 [0.81] 0.61 [0.80] 0.44 [0.72] 0.90 [0.95] 0.76 [0.88] 17.0
+Mistral AmbigQA 0.60 [0.79] 0.25 [0.62] 0.89 [0.95] 0.89 [0.95] 0.86 [0.93] 0.95 [0.98] 0.95 [0.98] 11.7
+HotpotQA 0.61 [0.80] 0.38 [0.67] 0.94 [0.97] 0.90 [0.95] 0.94 [0.97] 0.96 [0.98] 0.95 [0.98] 6.0
+NQ-Open 0.48 [0.73] 0.29 [0.64] 0.85 [0.93] 0.84 [0.92] 0.84 [0.92] 0.95 [0.98] 0.95 [0.98] 14.7
+TriviaQA 0.47 [0.72] 0.24 [0.61] 0.76 [0.88] 0.73 [0.86] 0.47 [0.74] 0.87 [0.94] 0.85 [0.93] 20.3
+Table 1: Cohen’s Kappa and [Macro-F1] scores showing the agreement of individual evaluators, Majority Vote (MV), and CLEV with human majority across models and tasks.
+Disagr.
+(%) indicates the percentage of disagreements in the majority voting (Llama+GPT+Mistral), i.e., when the third judge is called in CLEV.
+Correlation with human evaluation
+As depicted by consistently high Cohen’s kappa in Table 1, CLEV maintains strong agreement with human evaluation.
+This represents an improvement over individual model performance, where individual judges generally showed varying levels of agreement with human evaluation.
+LLM-based evaluators demonstrate strong abilities in recognizing semantic variations while maintaining the core meaning, especially when assessing responses that use different terminology or structural approaches to convey the same information.
+For instance, evaluators correctly identified that “Salma Hayek” and “Salma Hayek Pinault” refer to the same individual, acknowledging the semantic equivalence despite differences in phrasing.
+Similarly, when assessing responses that use different terms for the same entity, such as recognizing “Nick Fury, Agent of S.H.I.E.L.D.” as part of the broader “Marvel” universe, the evaluators effectively maintain the core meaning and contextual relevance.
+Their explanations show systematic assessment patterns that combine multiple evaluation criteria (e.g., factual accuracy and contextual relevance).
+CLEV reducing third-judge calls by 80–95% while maintaining the performance across tasks.
+As depicted in Table 1, even on TriviaQA, where disagreement is high, CLEV scores remain in substantial to perfect agreement.
+EM underestimates and BERTScore overestimates.
+Interestingly, EM typically accomplishes better correlation with human evaluation on the instance-level in Table 1 than neural-based BERTScore.
+EM’s strict nature leads to lower overall performance, but its high precision ensures that when it identifies a match, it strongly aligns with human annotations.
 In contrast, BERTScore takes a more lenient approach to semantic matching.
-Although this leniency produces higher raw scores, it introduces more false positives, consequently reducing instance-level agreement with human judgments.
-This pattern emerges clearly in many models and tasks such as when evaluating Llama-3.1-70B on AmbigQA, EM shows a raw score of 42.3% but achieves a Macro-F1 of 0.744, while BERTScore indicates a higher raw score of 63.0% but a lower Macro-F1 of 0.641.
-Table 2: Macro-F1 scores of various evaluators applied to different candidate LLMs and associated tasks.
-Higher scores indicate better performance.
-DAFE consistently achieves the highest Macro-F1 across all evaluated settings.
-Evaluators
-LLMs Tasks EM BS DeepSeek Llama GPT Mixtral Mistral DAFE
-Llama AmbigQA 0.744 0.641 0.948 0.944 0.922 0.912 0.929 0.955
-HotpotQA 0.778 0.745 0.942 0.939 0.949 0.910 0.916 0.976
-NQ-Open 0.653 0.718 0.898 0.916 0.896 0.907 0.869 0.964
-TriviaQA 0.612 0.782 0.726 0.772 0.717 0.695 0.640 0.842
-GPT AmbigQA 0.792 0.622 0.976 0.972 0.949 0.930 0.927 0.984
-HotpotQA 0.794 0.623 0.903 0.977 0.987 0.936 0.966 0.993
-NQ-Open 0.703 0.606 0.904 0.942 0.911 0.911 0.914 0.978
-TriviaQA 0.646 0.681 0.796 0.824 0.700 0.789 0.730 0.887
-Mixtral AmbigQA 0.760 0.666 0.948 0.948 0.891 0.955 0.944 0.975
-HotpotQA 0.761 0.657 0.960 0.970 0.966 0.930 0.970 0.987
-NQ-Open 0.650 0.649 0.912 0.939 0.863 0.950 0.908 0.956
-TriviaQA 0.625 0.695 0.829 0.812 0.803 0.838 0.716 0.882
-Mistral AmbigQA 0.792 0.622 0.947 0.947 0.947 0.947 0.930 0.977
-HotpotQA 0.796 0.673 0.951 0.969 0.951 0.947 0.969 0.979
-NQ-Open 0.726 0.639 0.898 0.925 0.919 0.939 0.920 0.976
-TriviaQA 0.718 0.608 0.925 0.879 0.863 0.822 0.735 0.927
-Figure 3: Heatmap illustrating the performance of four individual LLM judges on HotpotQA.
-Each cell value represents percentages (%).
-Rows represent predicted outcomes (P: Positive, N: Negative), while columns represent actual outcomes.
-See Appendix C for full results.
-4.2 Analysis
-In our experiments, candidate LLMs generated 7,500 outputs for the given tasks, with each evaluator producing 7,500 corresponding evaluations.
-We randomly sampled 100 error cases (50 false positives and 50 false negatives) from each evaluator to understand their behavior.
-Given EM had only 11 false positives, we included all of them in our analysis.
-Due to space constraints, we moved the detailed analysis of EM and BERTScore to Appendix C and focused exclusively on the LLM-as-a-judge method here.
-LLM-based evaluators demonstrate strong abilities in recognizing semantic variations while maintaining the core meaning, especially when assessing responses that use different terminology or structural approaches to convey the same information.
-For instance, in the evaluation examples, evaluators correctly identified that “Salma Hayek” and “Salma Hayek Pinault” refer to the same individual, acknowledging the semantic equivalence despite differences in phrasing.
-Similarly, when assessing responses that use different terms for the same entity, such as recognizing “Nick Fury, Agent of S.H.I.E.L.D.” as part of the broader “Marvel” universe, the evaluators effectively maintain the core meaning and contextual relevance.
-Their explanations show systematic assessment patterns that combine multiple evaluation criteria including factual accuracy, logical coherence, and contextual relevance.
-LLMs are prone to hallucination in justification (Zhang et al., 2023), where they fabricate reasoning to support their evaluations, produce detailed but incorrect explanations, or reference non-existent criteria or standards.
-In LLM judges, false positives and negatives (e.g., see Figure 3) often result from overlooking critical distinctions between candidate LLM outputs and failing to account for the specificity required by the reference answer.
+As a result, it shows more false positives, consequently reducing instance-level agreement with human judgments.
+4.1 Impact of Selective Third Judge
+A core benefit of CLEV is that it avoids unnecessary third-judge evaluations by only invoking a judge when the two primary judges disagree.
+Table 2 summarizes the frequency of such escalation events across five datasets and five candidate LLMs.
+Across all 7,500 evaluated instances (300 examples $\times$ 5 models $\times$ 5 tasks), only 17.6% required a third judge.
+This reduction is substantial when compared to the fixed three-judge majority voting setup, which always incurs 100% of third-judge calls.
+In other words, CLEV reduces evaluation overhead by approximately 80–95% while preserving agreement levels that match majority vote outcomes (see Table 1).
+Tasks with stable factual grounding (e.g., HotpotQA) trigger few escalations, whereas more temporally dynamic datasets like FreshQA show higher disagreement rates due to shifts in real-world facts.
+Tasks Third judge calls (of 1,500) Rate (%)
+AmbigQA 190 12.7
+FreshQA 543 36.2
+HotpotQA 120 8.0
+NQ-Open 212 14.1
+TriviaQA 253 16.9
+Total 1,318 / 7,500 17.6
+Table 2: Aggregated third-judge usage under CLEV across tasks.
+Each task has 1,500 evaluation instances per candidate model (300 examples $\times$ 5 candidate LLMs), and the reported counts sum over all five models  [^2] We also included DeepSeek as our fifth model in the Appendix.
+See Table 8..
+Lower values indicate fewer disagreement-triggered escalations.
+4.2 Error analysis
+We randomly sampled 100 error cases (50 false positives and 50 false negatives) from each evaluator to study their behavior. Given EM had 11 false positives, we included all of them in our analysis.
+LLMs are prone to hallucination in justification where they fabricate reasoning to support their evaluations, and produce detailed but incorrect explanations.
+In LLM judges, false positives and negatives often result from overlooking critical distinctions between candidate LLM outputs and failing to account for the specificity required by the reference answer.
 This pattern is particularly noticeable in Mistral 7B, where the model disregards the ground truth and provides evaluations influenced by unknown factors.
-For example, when evaluating candidate GPT-3.5’s response “The foreign minister of Germany who signed the Treaty of Versailles was Hermann Müller.” which is correct according to the reference answer “Hermann Müller” and human evaluation, Mistral 7B as-a-judge incorrectly marked this response as false and fabricated reasoning “Hermann Müller was the Chancellor of Germany, not the Foreign Minister.
-The Foreign Minister of Germany who signed the Treaty of Versailles was Gustav Stresemann.” in support of its decision.
-The same problem can also be attributed to inconsistent evaluations.
-Because when Mistral 7B acted as a candidate for the same question, its response to the question is completely different: “The Treaty of Versailles was signed by Matthias Erzberger, a German politician who served as the President of the German National Assembly at the time”.
-There are also alternative interpretations of this issue, such as ambiguity in the question, but we leave a deeper exploration of these aspects to future work.
-We observe a different pattern in some judges, specifically, GPT-3.5 and Mixtral 8x7B which focuses more on specificity.
+For example, when evaluating candidate GPT-3.5’s response “The foreign minister of Germany who signed the Treaty of Versailles was Hermann Müller.” which is correct according to the reference answer “Hermann Müller” and human evaluation, Mistral 7B as-a-judge incorrectly marked this response as false and fabricated reasoning in support of its decision.
+Specificity bias occur in some judges.
 This approach shifts the evaluation towards false negatives by missing semantically similar but structurally different answers.
 We found many cases when such evaluators failed to account for valid variations in phrasing or granularity, focusing instead on rigid adherence to the reference answer.
-Compounding these issues are reasoning errors within the evaluators’ own explanations, which often contain fabrications, circular logic, or overconfident assertions.
-By insisting on correctness derived strictly from the reference, evaluators disregard valid alternative perspectives and can even mischaracterize or invert the facts in their attempts to justify their decisions.
-This dynamic leaves little room for nuance or ambiguity, and it pushes the evaluation process away from fair, context-sensitive assessment toward rigid, and sometimes inaccurate, verdicts.
-Verbosity (Ye et al., 2024) emerges as a subtle source of bias, where more elaborate answers are sometimes overrated simply due to their detail and fluency, while concise yet correct responses are undervalued.
-This misplaced emphasis leads to irrelevant judgment criteria, such as praising the presence of irrelevant information or penalizing perfectly valid but succinct answers.
-We also found that LLM-based judges encounter challenges in multiple reference answers and more open-ended questions.
-This confusion is especially pronounced in the TriviaQA where the diversity and flexibility of valid responses present challenges for the judges’ ability to consistently recognize and evaluate a range of correct answers.
+Compounding these issues are reasoning errors within the evaluators’ own explanations, which often contain overconfident assertions.
 We found several temporal limitations in LLM-based evaluators.
-Although most of our datasets are older and the evaluator models are relatively up-to-date, we still observed instances where references to recent events, newly emerging terminology, or evolving contexts were misinterpreted.
-The FreshQA dataset (Vu et al., 2023), being recent, serves as a valuable testbed for assessing these temporal deficiencies.
-As shown in Table 3, LLM-based evaluators indicate deviation from human judgment on FreshQA compared to tasks that rely on older information, such as HotpotQA.
-Specifically, in dynamic or time-sensitive contexts, we found that LLM judges tend to hallucinate by consistently classifying candidate model responses as True, even when incorrect.
-For example, when presented with the question: “On what date did the Patriots last play the Miami Dolphins?” the LLM-generated response states: ‘‘The last time the New England Patriots played the Miami Dolphins was on January 1, 2023, during the NFL regular season.” Despite the correct reference answer being “November 24, 2024” the LLM evaluator not only failed to recognize the inaccuracy but also hallucinated an erroneous justification, stating: “The proposed answer correctly states the date the New England Patriots last played the Miami Dolphins as January 1, 2023, which matches the information provided.”
-Table 3: Performance (in Macro F1) of LLM judges on FreshQA.
+Despite using mostly older datasets with up-to-date models, we observed failures on recent events or evolving contexts.
+The FreshQA dataset (vu2023freshllms), being recent, highlights these temporal limitations.
+As shown in Table 3, LLM judges deviate more from human judgment on FreshQA compared to older datasets like HotpotQA.
+In dynamic or time-sensitive contexts, they often hallucinate by marking incorrect responses as True.
+For instance, when asked: “On what date did the Patriots last play the Miami Dolphins?”, the evaluator accepted an outdated answer, “January 1, 2023,” as correct, ignoring the reference “November 24, 2024,” and justifying its mistake.
+Because multiple judges rely on this same outdated parametric knowledge, FreshQA exhibits substantially higher disagreement between judges, who often converge on answers that were once correct but no longer reflect the current ground-truth.
 Evaluators
-LLMs DeepSeek Llama GPT Mixtral Mistral DAFE
-DeepSeek 0.714 0.692 0.715 0.614 0.724 0.830
-Llama 0.801 0.835 0.737 0.817 0.730 0.917
-GPT 0.659 0.695 0.824 0.780 0.746 0.891
-Mixtral 0.732 0.708 0.779 0.738 0.703 0.936
-Mistral 0.687 0.665 0.802 0.818 0.723 0.880
-Figure 4: Disagreement rates between primary judges (Llama+Mistral) across candidate LLMs and tasks.
-4.3 Disagreements between primary judges
-Figure 4 shows that disagreements between our primary judges, Llama-3.1 70B and Mistral 7B, mainly occur in the TriviaQA and FreshQA, with disagreement rates reaching 20.3% and 44.3%, respectively.
-Interestingly, higher disagreement rates between primary judges create a greater opportunity for DAFE to refine evaluations.
-As depicted in Figure 4, FreshQA (31.3% for Llama-70B, 39.7% for Mistral-7B) demonstrates the highest disagreement, allowing DAFE to improve Macro F1 scores (see Table 3).
-4.4 Impact of arbitration
-Our proposed arbitration approach significantly enhanced evaluation performance by resolving disputes through an independent judge, GPT-3.5-turbo (see Figure 5).
-Notably, in the AmbigQA, Macro F1 scores advanced from 72.9% to 86.6%, and Cohen’s Kappa increased from 0.467 to 0.773 (see Figure 7).
-These improvements highlight the pivotal role of the arbitrator in ensuring reliable and consistent evaluation outcomes.
-Figure 5: Comparison of Macro F1 scores before and after arbitration (see Appendix C for more results).
+Candid. Llama GPT Mistral CLEV Disagr. (%)
+Llama 0.835 0.737 0.730 0.917 31.3
+GPT 0.695 0.824 0.746 0.891 44.3
+Mixtral 0.708 0.779 0.703 0.936 37.3
+Mistral 0.665 0.802 0.723 0.880 39.7
+Table 3: Macro-F1 of judges on FreshQA and corresponding disagreement rates. Higher disagreement indicates a more frequent need for a resolving judge (CLEV).
 5 Related work
-Evaluation of natural language generation has traditionally relied on supervised signal-based metrics such as EM which evaluates the exact lexical match between generated outputs and reference answers.
-Despite its simplicity and efficiency, EM overlooks semantically equivalent variations, often penalizing accurate responses that use different phrasing (Wang et al., 2024a; Kamalloo et al., 2023).
-Other commonly used metrics including BLEU Papineni et al.
-(2002) and ROUGE (Lin, 2004) primarily focus on n-gram overlap with human written reference texts.
-Despite their widespread use, these metrics have significant limitations in capturing semantic subtleties and contextual relevance (Zhang et al., 2020).
-To address the limitations of conventional metrics, various model-based methods such as BERTScore Zhang et al.
-(2020) offer semantically informed evaluation.
-However, even BERTScore and similar embedding-based methods struggle to effectively evaluate open-ended generation Zheng et al.
-(2024); Sun et al.
-(2022).
-Recent advances in LLMs have unlocked new opportunities for automatic and context-aware evaluation (Li et al., 2024b; Chiang and Lee, 2023; Zheng et al., 2024).
-A key strength of LLM-based evaluators lies in their ability to operate in reference-free settings, where evaluation does not rely on pre-defined answers but instead leverages subjective criteria such as helpfulness, relevance, and coherence.
-This capability makes LLM evaluators particularly well-suited for assessing tasks where multiple valid responses exist or where human-like judgment is required (Li et al., 2024a).
-For instance, LLMs are frequently used in subjective evaluations such as pairwise comparison (“Which response is better?”) or single-response scoring (“How good is this response based on criteria X?” (Verga et al., 2024; Chan et al., 2024).
-LLM-based evaluators are specifically effective for tasks like summarization, where subjective criteria are central to evaluation (Liu et al., 2023).
-However, they are less effective for fact-based tasks such as free-form question-answering, where responses are either correct or incorrect and require explicit verification against reference answers.
-Furthermore, LLM-based evaluators face several challenges, particularly in ensuring consistency and fairness (Ye et al., 2024; Khan et al., 2024).
-In reference-free settings, the absence of a definitive ground truth increases the risk of bias in evaluations (Ye et al., 2024; Kim et al., 2024; Huang et al., 2024a).
-Common biases include positional bias, where LLMs may favor responses based on their order (Zheng et al., 2024; Khan et al., 2024), verbosity bias, which favors longer or more detailed responses (Huang et al., 2024b), and self-enhancement bias, where models may disproportionately prefer their own outputs (Zheng et al., 2024).
-These biases can distort evaluations and undermine the reliability of the results.
+Recent advances in LLMs have unlocked new opportunities for automatic and context-aware evaluation (li2024llmsasjudgescomp; chianglee2023large; 10.5555/3666122.3668142).
+Early work primarily focused on subjective evaluation tasks, such as pairwise comparison for ranking model outputs or single-response scoring for open-ended generation (chan2024chateval; 10.5555/3666122.3668142; chen2024humansllmsjudgestudy).
+These approaches leverage LLMs’ strong language understanding and reasoning capabilities to assess qualities like coherence, fluency, and helpfulness.
+However, such subjective criteria and continuous scoring schemes are less suitable for evaluating objective QA (badshah2025taletoolaugmented), where free-form answers are expected to be factually correct with respect to a reference answer (krumdick2025no; shi2024judgingjudg).
+To address this gap, recent work has explored reference-grounded LLM-as-judge approaches that instruct LLMs to directly verify candidate answers against gold references and output binary True/False verdicts (kamalloo-etal-2023-evaluating; NEURIPS2023_f323d594).
+However, this line of work also highlights key challenges: individual LLM judges can be inconsistent across questions, may over-rely on parametric knowledge when references are incomplete, and often require additional judges or majority voting to stabilize decisions (khan2024debating).
+Furthermore, single judges often lack robustness to adversarial inputs and may hallucinate justifications that appear plausible but are factually incorrect (hu2024llmbasedeval).
+To mitigate the reliability issues of single-judge systems, multi-judge setups have been explored (badshah-sajjad-2025-reference; verga2024replacing).
+These approaches employ multiple diverse LLM judges and aggregate their decisions, typically through majority voting, to improve reliability and alignment with human evaluation.
+badshah-sajjad-2025-reference demonstrated that using multiple models reduces individual model biases and improves agreement with human judgments.
+However, these approaches typically rely on fixed ensembles, where all judges are invoked for every instance, regardless of whether the initial judges already agree.
+This design leads to substantial and unnecessary computational overhead, especially when disagreements are rare.
+Our work builds on these foundations that invoke a third judge only when needed, achieving the reliability benefits of multi-judge voting while saving compute.
 6 Conclusion
-We present DAFE, a framework designed to evaluate free-form question-answering by leveraging LLMs.
-Our findings demonstrate that individual LLM judges are reliable alternatives to traditional lexical and neural-based metrics, offering closer alignment with human evaluations.
-However, relying solely on individual judges poses challenges including inherent biases and prompt sensitivity, which can affect evaluation performance.
-DAFE addresses these challenges through a dynamic arbitration mechanism.
-This design achieves near-perfect agreement with human evaluations, establishing DAFE as a trustworthy and reliable framework for evaluating open-ended language generation tasks.
-In the future, we aim to explore DAFE by excluding reference answers and integrating LLM agents with tools-interacting capabilities for evaluation.
+We studied the challenge of evaluating free-form QA, where traditional metrics fail to capture semantic correctness and single LLM judges may be inconsistent.
+To address this, we introduced CLEV, a lightweight multi-judge framework that preserves the reliability of majority voting while avoiding its computational cost.
+By invoking a third judge only when necessary, CLEV achieves substantial to near-perfect agreement with human evaluation while reducing redundant evaluation calls by 80–95% across tasks.
+This makes LLM-based evaluation more scalable for large-scale benchmarking.
+While our study focuses on binary correctness, future work can extend CLEV to capture dimensions such as partial correctness.
+Looking forward, an important direction is reference-free evaluation, where judges are supported by external verifiers (e.g., retrieval) to assess correctness in settings where explicit references are lacking.
 7 Limitations
-We acknowledge certain limitations in our study.
-The accuracy of evaluations depends on the quality and clarity of reference answers, which serve as the basis for determining correctness.
-Inconsistent or ambiguous references could affect evaluation outcomes.
-Similarly, this study primarily uses binary verdicts which might overlook detailed aspects of responses that could be captured through more comprehensive evaluation criteria.
-Furthermore, while we conducted an error analysis of LLM judges and automatic metrics, there may be error cases that were not identified during our manual review, leaving gaps in understanding the full spectrum of evaluation inaccuracies.
-Finally, our study focuses exclusively on English, and the applicability of our approach to other languages, particularly morphologically rich or resource-scarce ones, remains unexplored.
+We acknowledge certain limitations in our study: 1) The accuracy of evaluations depends on the quality and clarity of reference answers, which serve as the basis for determining correctness.
+Incorrect or ambiguous references could affect evaluation outcomes.
+2) While we conducted an error analysis of LLM judges and automatic metrics, there may be error cases that were not identified during our manual review, leaving gaps in understanding the full spectrum of evaluation inaccuracies.
+3) We acknowledge that our current study is limited to English QA datasets.
+Extending CLEV to multilingual settings is a promising and important future direction.
+4) Since the evaluation is binary, which is standard for free-form QA, signals around partial correctness, fluency, reasoning, and justification are not captured.
+To tackle this, we also collected detailed rationales that justify each verdict.
+In our analysis, we leveraged these rationales to capture various insights.
+We acknowledge the limitation of binary verdicts in representing partial correctness, but we emphasize that this design choice reduces the subjectivity inherent in continuous scoring and promotes clearer agreement between judges.
+5) CLEV assumes that primary judges maintain a minimum reliability threshold (based on $\kappa$ and Macro-F1) established during calibration.
+These thresholds require recalibration when adopting new judge models, domains, or languages.
+Thus, applying CLEV to new settings may involve additional upfront calibration cost.
+Acknowledgment
+We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC), Canada Foundation for Innovation (CFI), and Research Nova Scotia.
+Advanced computing resources are provided by ACENET, the regional partner in Atlantic Canada, and the Digital Research Alliance of Canada.
 Appendix A Free-form Question-Answering
-In our experiments, we include AmbigQA (Min et al., 2020), FreshQA (Vu et al., 2023), HotpotQA (Yang et al., 2018), Natural Questions (Kwiatkowski et al., 2019), and TriviaQA (Joshi et al., 2017).
+In our experiments, we include AmbigQA (minetal2020ambigqa), FreshQA (vu2023freshllms), HotpotQA (yang2018hotpotqa), Natural Questions (kwiatkowskietal2019), and TriviaQA (joshi2017triviaqa).
   • AmbigQA: Focuses on 14K ambiguous questions derived from NQ, requiring systems to identify multiple valid interpretations and generate disambiguated questions alongside corresponding answers.
@@
 We randomly sampled 300 examples from each dataset using Seed 42.
-Appendix B Human evaluation
+Appendix B Baselines
+Exact Match (EM):
+For our selected datasets and also free-form QA tasks, EM serves as a standard lexical matching metric to evaluate candidate LLM performance (izacardgrave2021; 105555; gou2024criticllms).
+Due to the verbose nature of LLM-generated responses, we adapt EM to classify an answer as correct if any golden answer $r_{i}\in R$ appears within the generated response $\bar{y}$ (i.e., $r_{i}\subseteq\bar{y}$), rather than requiring complete strict string equality (i.e., $\bar{y}=r_{i}$).
+BERTScore:
+We use BERTScore (ZhangKWWA20) which measures similarity by comparing contextualized word embeddings derived from a pre-trained BERT model.
+This enables the evaluation to focus on semantic correctness rather than exact lexical matches.
+As BERTScore is based on continuous values between -1 and 1, we set a threshold of $\tau=0.5$ to convert continuous similarity scores into binary 0 and 1.
+The purpose of this conversion is to allow direct comparison with other evaluation methods.
+For our implementation, we use the microsoft/deberta-xlarge-mnli [^3] https://huggingface.co/microsoft/deberta-xlarge-mnli model (he2021deberta).
+Majority voting
+This uses three fixed LLM judges to independently evaluate each instance.
+The final decision is determined by a simple majority across the three verdicts.
+Unlike CLEV, which selectively invokes the third judge only in cases of disagreement, this method uniformly engages all judges, leading to a higher computational cost.
+Appendix C Human evaluation
 This section provides detailed guidelines for human annotators responsible for evaluating the outputs of candidate LLMs.
-The goal is to ensure consistency and objectivity across all evaluations.
+The goal is to ensure consistency and objectivity across all evaluations (yu-etal-2024-latent).
 These guidelines provide clear instructions for assessing each model’s response based on its alignment with the reference answer and contextual relevance.
-B.1 Guidelines
+LLMs AmbigQA FreshQA HotpotQA NQ-Open TriviaQA
+DeepSeek 0.975 0.949 0.986 0.889 0.456 ($\kappa$ paradox)
+Llama 0.945 0.962 0.973 0.985 0.935
+GPT 0.989 0.973 0.982 0.990 0.948
+Mixtral 0.981 0.945 0.996 0.977 0.936
+Mistral 0.978 0.932 0.981 0.978 0.975
+Table 4: Fleiss’ Kappa scores of human annotators across models and tasks.
+LLMs AmbigQA FreshQA HotpotQA NQ-Open TriviaQA
+DeepSeek 99.0% 98.0% 99.7% 92.0% 90.0%
+Llama 96.3% 98.0% 98.0% 99.0% 99.0%
+GPT 99.3% 99.3% 98.7% 99.3% 99.0%
+Mixtral 98.7% 98.0% 99.7% 98.3% 98.3%
+Mistral 98.3% 97.0% 98.7% 98.3% 99.0%
+Table 5: Human annotators percent agreement scores across candidate models and tasks.
+C.1 Guidelines
 Dear Evaluator,
@@
 To maintain reliability, document any challenging cases requiring further discussion with other annotators.
-B.2 Inter human annotator agreement
-We calculate Fleiss’ Kappa ($\kappa$(Fleiss and Cohen, 1973) to assess inter-rater reliability among human annotators. Table 4 and 5 show the inter-annotator agreement across models and tasks.
-LLMs AmbigQA FreshQA HotpotQA NQ-Open TriviaQA
-DeepSeek 0.975 0.949 0.986 0.889 0.456 ($\kappa$ paradox)
-Llama 0.945 0.962 0.973 0.985 0.935
-GPT 0.989 0.973 0.982 0.990 0.948
-Mixtral 0.981 0.945 0.996 0.977 0.936
-Mistral 0.978 0.932 0.981 0.978 0.975
-Table 4: Fleiss’ Kappa scores of human annotators across models and tasks.
+Appendix D Evaluation Metrics
+We compute Fleiss’ Kappa ($\kappa$(fleiss1973equivalence) and percent agreement to assess inter-rater reliability among human annotators.
+Similarly, we use Cohen’s kappa (mchugh2012interrater) to find the agreement between each evaluator and the human majority to obtain instance-level comparison.
+Due to the high-class imbalance in TriviaQA, kappa scores can be misleadingly low despite high raw agreement - a known limitation called the “kappa paradox” (cicchetti1990high).
+Therefore, we treat the evaluation as a binary classification task where we consider each evaluator’s predictions against the human majority and report Macro-F1 scores which give equal weight to both classes regardless of their frequency in the selected random samples.
+To quantify the efficiency of our approach, we report the disagreement rate between the two primary judges that indicates how often the third model is required, thereby revealing the reduction in third-model usage compared to always employing three judges.
+Formally,
+$\text{Disagreement rate (\%)}=\left(\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}[V_{i_{1}}\neq V_{i_{2}}]\right)\times 100$
+where $N$ is the total number of evaluation instances and $\mathbb{I}[\cdot]$ is the indicator function that equals 1 when the condition is satisfied and 0 otherwise.
+D.1 Inter-human annotator agreement
+We calculate Fleiss’ Kappa ($\kappa$(fleiss1973equivalence) and percent agreement to assess inter-rater reliability among human annotators.
+Fleiss’ Kappa is defined as:
+$\kappa=\frac{\bar{P}-P_{e}}{1-P_{e}},$
+where $\bar{P}$ is the average observed agreement among annotators, and $P_{e}$ is the expected agreement by chance.
+Percent agreement is calculated as:
+$\text{Percent Agreement}=\left(\frac{\text{Agreements}}{\text{Total Annotations}}\right)\times 100$
+Table 4 and 5 show the inter-annotator agreement across models and tasks.
 The results demonstrate high reliability, with Fleiss’ Kappa scores consistently above 0.93 for most tasks.
@@
 This indicates a variance in human ratings for these tasks.
-LLMs AmbigQA FreshQA HotpotQA NQ-Open TriviaQA
-DeepSeek 99.0% 98.0% 99.7% 92.0% 90.0%
-Llama 96.3% 98.0% 98.0% 99.0% 99.0%
-GPT 99.3% 99.3% 98.7% 99.3% 99.0%
-Mixtral 98.7% 98.0% 99.7% 98.3% 98.3%
-Mistral 98.3% 97.0% 98.7% 98.3% 99.0%
-Table 5: Human annotators percent agreement scores across candidate models and tasks.
-Appendix C Additional results
+Appendix E Additional results
 This section provides further results and analysis of conventional metrics and LLM-based evaluators.
-Table 6 illustrates the overall performance of candidate LLMs obtained through various evaluators.
-Unlike lexical matching and neural-based metrics, each LLM-as-a-judge indicates overall performance close to the human majority.
-Automatic metrics like EM severely underestimate the candidate LLMs’ performance.
-On the other hand, BERTScore tends to overestimate the performance.
+As evidenced by high Cohen’s kappa and Macro F1 scores in Table 6 and 7, CLEV maintains a strong alignment with human evaluation.
+This represents a substantial improvement over individual model performance, where individual judges generally showed varying levels of agreement with human evaluation.
+Overall, LLM-as-a-judge works better with larger models.
+This is particularly noticeable in Llama and GPT, which achieve greater performance across AmbigQA, HotpotQA, and NQ-Open compared to smaller models.
+This indicates an important scaling law in evaluation capability (kaplan2020scalinglawsneurallanguage; 10.5555/3666122.3668142; openai2024gpt4technicalreport).
+However, we also found that the most advanced models are not always guaranteed to be the best evaluators.
+We observed slightly comparable performance through the small open-source Mistral7B.
+For instance, when evaluating candidate Mixtral 8x7B on AmbigQA (see Table  7), Mistral 7B as-a-judge outperformed (0.944) judge GPT-3.5-turbo (0.891).
+Regardless, we observe relatively lower Macro-F1 scores for all LLM judges in TriviaQA.
+Evaluators
+LLMs Tasks EM BS Llama GPT Mixtral Mistral CLEV
+Llama AmbigQA 0.518 0.283 0.888 0.844 0.824 0.858 0.911
+HotpotQA 0.577 0.498 0.877 0.899 0.820 0.832 0.953
+NQ-Open 0.381 0.437 0.833 0.793 0.816 0.738 0.927
+TriviaQA 0.281 0.564 0.547 0.439 0.396 0.299 0.684
+GPT AmbigQA 0.561 0.252 0.944 0.897 0.861 0.853 0.967
+HotpotQA 0.604 0.300 0.953 0.973 0.873 0.933 0.987
+NQ-Open 0.453 0.218 0.884 0.824 0.824 0.829 0.956
+TriviaQA 0.335 0.364 0.650 0.401 0.580 0.467 0.775
+Mixtral AmbigQA 0.546 0.337 0.896 0.781 0.909 0.887 0.951
+HotpotQA 0.546 0.349 0.940 0.933 0.859 0.940 0.973
+NQ-Open 0.371 0.301 0.879 0.728 0.899 0.815 0.913
+TriviaQA 0.317 0.390 0.625 0.605 0.678 0.436 0.764
+Mistral AmbigQA 0.599 0.254 0.893 0.893 0.893 0.860 0.953
+HotpotQA 0.605 0.383 0.937 0.902 0.895 0.937 0.958
+NQ-Open 0.484 0.291 0.851 0.838 0.878 0.840 0.953
+TriviaQA 0.467 0.239 0.758 0.725 0.645 0.470 0.854
+Table 6: Cohen’s Kappa scores displaying the agreement levels of individual and multiple (CLEV) evaluators with human judgments across candidate models and tasks.
+Evaluators
+LLMs Tasks EM BS Llama GPT Mixtral Mistral CLEV
+Llama AmbigQA 0.744 0.641 0.944 0.922 0.912 0.929 0.955
+HotpotQA 0.778 0.745 0.939 0.949 0.910 0.916 0.976
+NQ-Open 0.653 0.718 0.916 0.896 0.907 0.869 0.964
+TriviaQA 0.612 0.782 0.772 0.717 0.695 0.640 0.842
+GPT AmbigQA 0.792 0.622 0.972 0.949 0.930 0.927 0.984
+HotpotQA 0.794 0.623 0.977 0.987 0.936 0.966 0.993
+NQ-Open 0.703 0.606 0.942 0.911 0.911 0.914 0.978
+TriviaQA 0.646 0.681 0.824 0.700 0.789 0.730 0.887
+Mixtral AmbigQA 0.760 0.666 0.948 0.891 0.955 0.944 0.975
+HotpotQA 0.761 0.657 0.970 0.966 0.930 0.970 0.987
+NQ-Open 0.650 0.649 0.939 0.863 0.950 0.908 0.956
+TriviaQA 0.625 0.695 0.812 0.803 0.838 0.716 0.882
+Mistral AmbigQA 0.792 0.622 0.947 0.947 0.947 0.930 0.977
+HotpotQA 0.796 0.673 0.969 0.951 0.947 0.969 0.979
+NQ-Open 0.726 0.639 0.925 0.919 0.939 0.920 0.976
+TriviaQA 0.718 0.608 0.879 0.863 0.822 0.735 0.927
+Table 7: Macro-F1 scores of individual and multiple (CLEV) evaluators applied to different candidate LLMs and associated tasks.
+Unlike lexical matching and neural-based metrics, each LLM-as-a-judge shows overall performance close to the human majority.
+The proposed CLEV method consistently achieves comparable or slightly better alignment with the human majority.
+Conventional metrics such as EM severely underestimate the candidate LLMs’ performance.
+Contrarily, BERTScore tends to overestimate the performance except in some cases such as when evaluating Llama on AmbigQA and NQ-Open.
 EM underestimates performance because it requires a candidate’s response to exactly match one of the reference answers.
@@
 For example, in cases where minor differences in wording (e.g., “The Treaty of Versailles was signed in 1919.” versus “The Treaty of Versailles ended in 1919.”) lead to opposing factual claims, BERTScore still scores the response high due to its emphasis on matching tokens (e.g., “signed” versus “ended”).
-Additionally, verbosity bias and threshold instability—where a default threshold (threshold = 0.5) is arbitrarily set—further inflate its raw accuracy.
+Additionally, verbosity bias and threshold instability further inflate its raw accuracy (see Table 10).
 However, when comparing raw accuracy with instance-level agreement metrics like Cohen’s kappa, which adjusts for class imbalance and penalizes asymmetric errors, the limitations of BERTScore become apparent.
-LLMs Tasks Evaluators
-EM BS HM DeepSeek Llama GPT Mixtral Mistral
-DeepSeek AmbigQA 56.3 80.0 84.3 86.3 73.7 75.0 62.3 93.3
-FreshQA 31.3 88.0 84.3 84.7 82.7 75.3 58.0 82.3
-HotpotQA 38.6 78.4 57.7 58.0 51.0 51.0 52.7 57.7
-NQ-Open 35.0 78.3 60.3 64.7 63.7 61.3 55.3 68.3
-TriviaQA 77.3 90.7 94.3 90.7 94.0 91.7 81.7 89.7
-Llama AmbigQA 42.3 63.0 67.0 64.0 65.3 64.7 63.0 66.0
-FreshQA 25.6 81.3 77.7 81.3 78.3 72.7 71.0 62.3
-HotpotQA 34.3 67.7 56.3 56.7 58.3 54.0 50.7 52.7
-NQ-Open 31.7 61.7 66.3 62.3 62.7 60.0 59.0 66.7
-TriviaQA 74.3 94.0 94.7 88.0 90.3 90.0 88.7 84.7
-GPT AmbigQA 49.7 78.0 71.7 70.3 70.0 68.0 65.7 71.0
-FreshQA 24.6 89.3 70.7 58.0 51.7 78.7 83.0 83.3
-HotpotQA 33.7 80.0 54.0 50.3 53.0 52.7 51.7 54.0
-NQ-Open 36.3 74.0 65.3 65.3 62.7 59.0 59.0 67.0
-TriviaQA 74.3 95.3 93.0 90.0 89.3 90.7 89.7 86.3
-Mixtral AmbigQA 37.7 70.3 61.7 58.7 57.3 62.0 59.3 61.7
-FreshQA 18.6 89.7 86.0 72.3 67.0 87.0 85.0 77.7
-HotpotQA 25.0 69.7 47.0 46.3 45.3 45.7 44.7 46.0
-NQ-Open 23.7 63.7 56.7 54.0 52.7 47.7 52.3 59.7
-TriviaQA 64.7 91.3 90.7 83.7 86.3 89.7 86.0 85.3
-Mistral AmbigQA 31.0 61.7 49.7 47.7 46.3 47.7 46.3 53.3
-FreshQA 15.6 80.0 81.7 60.7 59.0 83.7 84.0 86.0
-HotpotQA 23.7 64.7 40.0 39.3 39.0 38.0 37.0 39.0
-NQ-Open 22.7 60.0 46.0 41.3 40.0 43.3 41.3 50.0
-TriviaQA 62.0 94.3 83.7 78.0 81.3 81.0 79.7 85.0
-Table 6: Raw performance of candidate LLMs across free-form QA tasks evaluated through various methods. HM represents Human Majority and BS denotes BERTScore.
-C.1 Impact of arbitration on dispute resolution
-Figure 6 illustrates the impact of arbitration on resolving disagreements between primary judges.
-Arbitration, facilitated by GPT-3.5 as the tiebreaker, consistently improves performance across all tasks, particularly in FreshQA and TriviaQA, where Macro F1 increases by up to 21.5 points.
-In contrast, tasks like AmbigQA and HotpotQA, where primary judges initially exhibit stronger agreement, show smaller but still meaningful improvements.
-This highlights the critical role of arbitration in enhancing agreement and achieving closer alignment with ground truth, especially in cases of significant disagreement among primary judges.
-Notably, evaluations of DeepSeek-v3 exhibit higher disagreement between Llama-3.1-70B and Mistral-7B, particularly in FreshQA (28.3%) and AmbigQA (25.7%).
-From our analysis, we did not find strong evidence explaining why DeepSeek-v3 leads to higher disagreement between the primary judges.
-Figure 6: Impact of arbitration on disagreements between primary judges.
-Note that we used Llama-3.1-70B and Mistra 7B as primary judges.
-GPT-3.5-turbo is only utilized when disagreements are found.
-The models given in the figure are candidate LLMs which generate outputs for the given tasks and are then evaluated through DAFE.
-We observed substantial enhancements in Cohen’s Kappa scores across several tasks.
-For instance, as illustrated in Figure 7, in the AmbigQA Cohen’s Kappa increased from 0.881 to 0.911 for Llama.
-Similarly, in the same task, Cohen’s Kappa from 0.467 to 0.773 for candidate DeepSeek.
-These improvements demonstrate that the arbitration mechanism effectively enhances the reliability and consistency of evaluations, particularly in complex and ambiguous tasks where primary judges are more likely to disagree.
-Some Cohen’s Kappa scores remain relatively low, particularly in FreshQA and DeepSeek-evaluated outputs.
-This is partially explained by the Kappa Paradox, where high agreement on extreme cases (e.g., clear correct/incorrect responses) and unbalanced class distributions can artificially lower the Kappa scores.
-In such cases, even when evaluators mostly agree, Cohen’s Kappa can appear lower than expected.
-Despite this, the arbitration process effectively mitigates inconsistencies, especially in tasks involving evolving knowledge and nuanced interpretations, such as FreshQA.
-Figure 7: Comparison of Cohen’s kappa scores before and after arbitration (GPT-3.5-turbo as arbitrator). The performance is illustrated across candidate LLMs and tasks.
-C.2 Cost analysis
-Human evaluation is the gold standard for assessing LLM-generated responses, but it is expensive and time-consuming.
-In our setup, we employed three human annotators who volunteered their efforts.
-However, if these annotators were compensated based on standard annotation rates, the cost of evaluating such outputs would be significantly higher.
-On the other hand, GPT-3.5-turbo, acting as an arbitrator in DAFE, incurs a cost that depends on the number of arbitration cases.
-In our evaluations, GPT-3.5-turbo was invoked 1,318 times, with an estimated total cost of \$0.59, which increases to \$5.40 if a 2048 max token setting is used (see Table 7).
-Since GPT-3.5 is only invoked when primary judges disagree, this selective arbitration substantially reduces overall evaluation expenses while maintaining high reliability in assessments.
-Rather than relying on a single model for evaluation, this multi-model arbitration approach enhances trust by mitigating biases and weaknesses inherent in any individual model.
-By invoking the arbitrator only when disagreements occur (rather than evaluating all responses), DAFE reduces arbitration usage by 82–94% compared to a majority-voting system. This leads to:
-  • Over 90% fewer third-judge inferences, drastically lowering computational demand.
-  • Up to 95% cost savings by avoiding redundant model evaluations.
-  • Better scalability, making it practical for large-scale deployments
-Table 7: Cost-efficiency analysis of DAFE: Summary of disagreement rates and tiebreaker usage across candidate models and tasks
-Candidate LLMs Tasks Samples Disagreement Rates (%) Tiebreaker Usage
+E.1 Cost analysis
+To assess the efficiency of CLEV, we track the number of times the third judge is invoked, which directly corresponds to disagreement between the two primary judges.
+As shown in Table 8, across 7,500 evaluation instances, the third judge is required only 1,318 times, representing just 17.6% of the total cases.
+This implies an 82.4% reduction in third-judge usage compared to a full majority-voting setup, where every instance would involve all three models.
+Disagreement rates vary across tasks and models.
+For example, GPT shows only 5.7% disagreement on HotpotQA, while FreshQA exhibits higher disagreement (up to 44.3%) for some judge combinations.
+This behavior allows CLEV to scale efficiently: it concentrates computational effort only where model uncertainty exists, minimizing redundant inference.
+In contrast to fixed-cost evaluation schemes, CLEV offers a cost-efficient alternative that maintains high evaluation quality while significantly reducing compute usage.
+Candidate LLMs Tasks Samples Disagreement Rates (%) Third Judge Usage
 DeepSeek AmbigQA 300 25.7 77
@@
 Total 7500 1318
-C.3 DeepSeek as the arbitrator
-To assess the impact of using DeepSeek as the arbitrator in DAFE, we conducted experiments by replacing GPT-3.5-turbo with DeepSeek.
-We evaluated this setup using different candidate models across multiple tasks.
-Specifically, we tested GPT-3.5 on TriviaQA, DeepSeek on NQ-Open, and Llama on FreshQA.
-The primary judges remained Llama and Mistral, and arbitration was invoked only in cases of disagreement.
-Our findings indicate that DeepSeek as the arbitrator achieves strong performance, with Macro-F1 scores of 91.23 on TriviaQA, 79.11 on NQ-Open, and 0.914 on FreshQA.
-C.4 Evaluating with one strong LLM-as-a-judge
+Table 8: Cost-efficiency analysis of CLEV: Summary of disagreement rates and third judge usage across candidate models and tasks
+E.2 DeepSeek as the third judge
+To assess the impact of using DeepSeek as the third judge in CLEV, we conducted experiments by replacing GPT-3.5-turbo with DeepSeek-R1 (deepseekai2025deepseekr1).
+We evaluated this setup using candidate GPT-3.5 on TriviaQA, candidate DeepSeek on NQ-Open, and candidate Llama on FreshQA.
+The primary judges remained Llama and Mistral, and the third judge (i.e., DeepSeek-R1) is invoked only in cases of disagreement.
+Our findings indicate that DeepSeek, as the third judge, achieves strong performance, with Macro-F1 scores of 91.23 on TriviaQA, 79.11 on NQ-Open, and 0.914 on FreshQA.
+E.3 Evaluating with GPT-4o as-a-judge
 While a single state-of-the-art evaluator can achieve strong performance in many cases, the dual-LLM framework remains critical for ensuring robustness, particularly in high-stakes or ambiguous scenarios.
@@
 With this configuration, GPT-4o as the evaluator achieved a Macro-F1 score of 0.946 on HotpotQA, demonstrating its exceptional capability.
-However, the same GPT-4o judge achieved only 0.784 on TriviaQA, which falls short of DAFE’s performance of 0.887.
+However, the same GPT-4o judge achieved only 0.784 on TriviaQA, which falls short of CLEV’s performance of 0.887.
 This shows that even the most advanced models show inconsistencies when evaluating free-form QA.
 This is particularly critical in precision-sensitive domains where minor errors can have outsized consequences.
-In such settings, DAFE’s ensemble approach acts as a safeguard.
-When employing DAFE with GPT-3.5-turbo as the arbitrator, we achieved an even higher Macro-F1 of 0.984 on HotpotQA, surpassing the performance of a single GPT-4o.
-Interestingly, when we experimented with DeepSeek as the arbitrator in DAFE, performance remained strong at 0.963 Macro-F1, indicating that DAFE’s benefits are not solely tied to a specific arbitrator model.
-C.5 Majority voting-based evaluation
-We conducted additional experiments utilizing a traditional majority voting approach for evaluating candidate LLM performance.
-In this setup, we employed three LLM judges of equal weight: Llama, GPT-3.5, and Mistral to evaluate candidate models generated response.
+In such settings, CLEV’s ensemble approach acts as a safeguard.
+When employing CLEV with GPT-3.5-turbo as the third judge, we achieved an even higher Macro-F1 of 0.984 on HotpotQA, surpassing the performance of a single GPT-4o.
+Interestingly, when we experimented with DeepSeek as the third judge in CLEV, performance remained strong at 0.963 Macro-F1, indicating that CLEV’s benefits are not solely tied to a specific third judge model.
+E.4 Majority voting-based evaluation
+We conducted additional experiments utilizing a traditional majority voting approach for evaluating candidate LLMs performance.
+Given $n$ annotators and a binary classification, the majority label is defined as:
+$y_{\text{majority}}=\begin{cases}1&\text{if }\sum_{i=1}^{n}y_{i}>\frac{n}{2},\\ 0&\text{otherwise},\end{cases}$
+where $y_{i}$ represents the label assigned by the $i$th annotator.
+In this setup, we employed three LLM judges of equal weight: Llama, GPT-3.5, and Mistral to evaluate candidate models’ generated responses.
 For every evaluation instance, each judge provided an independent binary verdict (True or False).
 The final decision is determined through a simple majority vote across these three verdicts.
-As presented in Table 8, DAFE matches or closely approaches the Macro F1 and Cohen’s Kappa scores of the three-judge majority across almost all tasks and candidate LLMs.
-For example, on HotpotQA, evaluating candidate Llama with DAFE achieves a Macro F1 of 97.6% (compared to 97.6% for majority voting) and a Cohen’s Kappa of 0.95, while for GPT-3.5 on AmbigQA, DAFE reaches a Macro F1 of 98.4% (versus 98.3% for majority voting), indicating a negligible performance difference.
-Even in high-disagreement tasks like TriviaQA, where the primary judges (e.g., Mistral) disagree 20.3% of the time, DAFE retains strong alignment (with a Macro F1 of 92.7 compared to 93.5 for majority voting).
-Minor deviations, such as the one observed for candidate Mixtral on TriviaQA (DAFE’s Macro F1 = 0.88 vs.
-0.95 for majority voting), reflect rare instances where both the primary judges and the arbitrator make errors, yet these outliers are substantially outweighed by the computational savings offered by selective arbitration.
-Table 8: Comparison between Majority Voting (Llama+GPT-3.5+Mistral) and DAFE (GPT-3.5 as arbitrator).
-For each candidate LLM and task, the table reports Macro F1 and Cohen’s Kappa scores under Majority Voting, the disagreement rate (in %), and the corresponding scores using DAFE.
-Candidate LLM Task Majority Voting Disagreement (%) DAFE
+As presented in Table 9, CLEV matches or closely approaches the Macro F1 and Cohen’s Kappa scores of the three-judge majority across almost all tasks and candidate LLMs.
+For example, on HotpotQA, evaluating candidate Llama with CLEV achieves a Macro F1 of 97.6% (compared to 97.6% for majority voting) and a Cohen’s Kappa of 0.95, while for GPT-3.5 on AmbigQA, CLEV reaches a Macro F1 of 98.4% (versus 98.3% for majority voting), indicating a negligible performance difference.
+Even in high-disagreement tasks like TriviaQA, where the primary judges (e.g., Mistral) disagree 20.3% of the time, CLEV retains strong alignment (with a Macro F1 of 92.7 compared to 93.5 for majority voting).
+Minor deviations, such as the one observed for candidate Mixtral on TriviaQA (CLEV’s Macro F1 = 0.88 vs.
+0.95 for majority voting), reflect rare instances where both the primary judges and the third judge make errors, yet these outliers are substantially outweighed by the computational savings offered by the selective third judge.
+Candidate LLM Task Majority Voting Disagreement (%) CLEV
 Macro F1 Kappa Macro F1 Kappa
@@
 TriviaQA 93.5 0.87 20.3 92.7 0.85
-C.6 Impact of prompt variations
+Table 9: Comparison between Majority Voting (Llama+GPT-3.5+Mistral) and CLEV (GPT-3.5 as the third judge).
+For each candidate LLM and task, the table reports Macro F1 and Cohen’s Kappa scores under Majority Voting, the disagreement rate (in %), and the corresponding scores using CLEV.
+Small differences in performance between the majority voting setup and CLEV are due to randomness (see Section E.5.1).
+E.5 Impact of prompt variations
 The effectiveness and consistency of LLM-based evaluation are significantly influenced by prompt design.
@@
 To analyze the robustness of the LLM judges in free-form QA, we conducted ablation studies on different prompt variations using Mistral as the candidate model and GPT as the judge.
-C.6.1 Consistency in judgment across multiple trials
+E.5.1 Consistency in judgment across multiple trials
 LLMs generate random text even at a temperature of 0.
-To assess whether this affects evaluation consistency, we repeated the same evaluation task five times for 100 Mistral-generated responses for HotpotQA.
-  • Verdict stability: GPT produced identical True/False verdicts in 100% of cases. This suggest that its binary decision-making process remains stable even across multiple trials.
+To assess whether this affects evaluation consistency, we repeated the same evaluation task five times for 300 Mistral-generated responses for HotpotQA.
+In this ablation study, we use GPT-3.5 as a judge.
+  • Verdict stability: GPT produced identical True/False verdicts in 98% of cases, with minor variations observed in only 1–2 instances.
+This suggest that its binary decision-making process remains largely stable even across multiple trials.
   • Explanation variability: While verdicts remained consistent, the rationales and explanations provided by GPT across trials, often cited different supporting facts for the same judgment.
-C.6.2 Few-shot vs. zero-shot prompting
+E.5.2 Few-shot vs. zero-shot prompting
 We investigated the impact of few-shot prompting where we included three task-specific examples in the prompt to guide the judge’s decision-making process.
@@
 For instance, multi-hop reasoning cases from HotpotQA, the judge model consistently followed the structure of the provided examples, even when the correct reasoning required a different approach.
-C.6.3 Explanation requirement: Binary verdict vs. justification-based evaluation
+E.5.3 Explanation requirement: Binary verdict vs. justification-based evaluation
 To test whether requiring the model to generate explanations alongside verdicts improves judgment reliability, we compared two settings:
@@
   • Reduced alignment with human judgment: Cohen’s Kappa agreement with human annotators dropped from 0.95 to 0.72, highlighting that rationale-based prompts lead to more stable and accurate decisions.
-C.6.4 Reason-first vs. verdict-first prompting
+E.5.4 Reason-first vs. verdict-first prompting
 In the verdict-first approach, the model is instructed to provide a True/False answer before justifying its decision, whereas in the reason-first approach, the model is asked to generate reasoning first and then conclude with a verdict.
 Experimental results showed no significant difference in accuracy or agreement scores between these two formats.
-C.7 G-Eval: reference-free evaluation of free-form question-answering
-Existing LLM-based evaluators such as G-Eval (Liu et al., 2023) are designed for reference-free, subjective tasks (e.g., summarization, dialogue), where evaluation criteria (e.g., coherence, fluency) are inherently ambiguous and scored on Likert scales.
+E.6 G-Eval: reference-free evaluation of free-form question-answering
+Existing LLM-based evaluators such as G-Eval (liuetal2023g) are designed for reference-free, subjective tasks (e.g., summarization, dialogue), where evaluation criteria (e.g., coherence, fluency) are inherently ambiguous and scored on Likert scales.
 These frameworks prioritize qualitative judgments rather than binary factual correctness.
-In contrast, DAFE is explicitly tailored for reference-dependent, objective evaluation in free-form QA, where answers are either factually correct or incorrect based on alignment with explicit ground-truth references.
-To validate this distinction, we tailored G-eval based method to investigate the capability of LLM-as-a-judge in reference-free settings.
-In this setting, we modify the evaluation prompt by excluding the reference answer $r$ and directly prompted the evaluator model as $P=\{x,\bar{y}\}$ along with instructions such as correctness.
+In contrast, CLEV is explicitly tailored for reference-dependent, objective evaluation in free-form QA, where answers are either factually correct or incorrect based on alignment with explicit ground-truth references.
+To validate this distinction, we tailored the G-eval (liuetal2023g) to investigate the capability of LLM-as-a-judge in reference-free settings.
+In this setting, we modify the evaluation prompt by excluding the reference answer $r$ and directly prompt the evaluator model as $P=\{x,\bar{y}\}$ along with instructions such as correctness.
 The performance of LLM-as-a-judge drastically changes in reference-free settings.
-Without access to the ground truth references, we observe a stark decline in evaluation capability across all models (see Table 9 and 10 values in blue).
+Without access to the ground truth references, we observe a stark decline in evaluation capability across all models (see Table 10 and 11 values in blue).
 This systematic deterioration spans all tasks and model combinations, though its severity varies by context.
-HotpotQA, with its demands for complex reasoning, exemplifies this challenge most clearly.
+HotpotQA and NQ-Open, with their demands for complex reasoning, exemplify this challenge most clearly.
 The substantial gap between reference-based and reference-free evaluation underscores the crucial role of reference answers in reliable assessment.
@@
 TriviaQA 62.0 94.3 83.7 81.3 [80.7] 81.0 [81.0] 79.7 [80.7] 85.0 [84.7]
-Table 9: Overall performance (Accuracy) of candidate LLMs across free-form QA tasks. Values [in blue] represent LLM-as-a-judge in the reference-free mood.
+Table 10: Overall performance (Raw Accuracy) of candidate LLMs across free-form QA tasks. Values [in blue] represent LLM-as-a-judge in the reference-free mood.
 Candidate LLMs Tasks Evaluators
-EM BERTScore Llama-3.1-70B GPT-3.5-turbo Mixtral-8x7B Mistral-7B
-Llama-3.1-70B AmbigQA 0.744 0.641 0.944 [0.629] 0.922 [0.604] 0.912 [0.669] 0.929 [0.631]
-HotpotQA 0.778 0.745 0.939 [0.628] 0.949 [0.574] 0.910 [0.665] 0.916 [0.640]
-NQ-Open 0.653 0.718 0.916 [0.606] 0.896 [0.560] 0.907 [0.639] 0.869 [0.622]
-TriviaQA 0.612 0.782 0.772 [0.772] 0.717 [0.628] 0.695 [0.678] 0.640 [0.633]
-GPT-3.5 AmbigQA 0.792 0.622 0.972 [0.686] 0.949 [0.603] 0.930 [0.596] 0.927 [0.553]
-HotpotQA 0.794 0.623 0.977 [0.566] 0.987 [0.521] 0.936 [0.543] 0.966 [0.494]
-NQ-Open 0.703 0.606 0.942 [0.671] 0.911 [0.544] 0.911 [0.601] 0.914 [0.536]
-TriviaQA 0.646 0.681 0.824 [0.817] 0.700 [0.690] 0.789 [0.760] 0.730 [0.701]
-Mixtral-8x7B AmbigQA 0.760 0.666 0.948 [0.704] 0.891 [0.636] 0.955 [0.654] 0.944 [0.622]
-HotpotQA 0.761 0.657 0.970 [0.587] 0.966 [0.470] 0.930 [0.582] 0.970 [0.577]
-NQ-Open 0.650 0.649 0.939 [0.652] 0.863 [0.517] 0.950 [0.590] 0.908 [0.529]
-TriviaQA 0.625 0.695 0.812 [0.800] 0.803 [0.754] 0.838 [0.818] 0.716 [0.725]
-Mistral-7B AmbigQA 0.792 0.622 0.947 [0.730] 0.947 [0.627] 0.947 [0.628] 0.930 [0.523]
-HotpotQA 0.796 0.673 0.969 [0.649] 0.951 [0.478] 0.947 [0.680] 0.969 [0.578]
-NQ-Open 0.726 0.639 0.925 [0.652] 0.919 [0.515] 0.939 [0.597] 0.920 [0.433]
-TriviaQA 0.718 0.608 0.879 [0.881] 0.863 [0.840] 0.822 [0.846] 0.735 [0.744]
-Table 10: Performance (Macro F1) of various evaluators across candidate LLMs and tasks. Values [in blue] represent LLM-as-a-judge in the reference-free mode.
-C.8 DAFE in multi-reference answers
-DAFE explicitly accommodates multiple gold reference answers by incorporating all available references into the judge LLM’s prompt during evaluation.
-For datasets like AmbigQA and TriviaQA, where questions often have multiple valid answers (e.g., synonyms, rephrased answers, or alternative factual representations), DAFE aggregates all reference answers into the judge’s input prompt (e.g., concatenating them as a comma-separated list).
+EM BERTScore Llama-3.1-70B GPT-3.5-turbo Mixtral-8x7B Mistral-7B CLEV
+Llama-3.1-70B AmbigQA 0.744 0.641 0.944 [0.629] 0.922 [0.604] 0.912 [0.669] 0.929 [0.631] 0.955 [0.637]
+HotpotQA 0.778 0.745 0.939 [0.628] 0.949 [0.574] 0.910 [0.665] 0.916 [0.640] 0.976 [0.623]
+NQ-Open 0.653 0.718 0.916 [0.606] 0.896 [0.560] 0.907 [0.639] 0.869 [0.622] 0.964 [0.610]
+TriviaQA 0.612 0.782 0.772 [0.772] 0.717 [0.628] 0.695 [0.678] 0.640 [0.633] 0.842 [0.747]
+GPT-3.5 AmbigQA 0.792 0.622 0.972 [0.686] 0.949 [0.603] 0.930 [0.596] 0.927 [0.553] 0.984 [0.607]
+HotpotQA 0.794 0.623 0.977 [0.566] 0.987 [0.521] 0.936 [0.543] 0.966 [0.494] 0.993 [0.522]
+NQ-Open 0.703 0.606 0.942 [0.671] 0.911 [0.544] 0.911 [0.601] 0.914 [0.536] 0.978 [0.575]
+TriviaQA 0.646 0.681 0.824 [0.817] 0.700 [0.690] 0.789 [0.760] 0.730 [0.701] 0.887 [0.882]
+Mixtral-8x7B AmbigQA 0.760 0.666 0.948 [0.704] 0.891 [0.636] 0.955 [0.654] 0.944 [0.622] 0.975 [0.650]
+HotpotQA 0.761 0.657 0.970 [0.587] 0.966 [0.470] 0.930 [0.582] 0.970 [0.577] 0.987 [0.536]
+NQ-Open 0.650 0.649 0.939 [0.652] 0.863 [0.517] 0.950 [0.590] 0.908 [0.529] 0.956 [0.563]
+TriviaQA 0.625 0.695 0.812 [0.800] 0.803 [0.754] 0.838 [0.818] 0.716 [0.725] 0.882 [0.858]
+Mistral-7B AmbigQA 0.792 0.622 0.947 [0.730] 0.947 [0.627] 0.947 [0.628] 0.930 [0.523] 0.977 [0.647]
+HotpotQA 0.796 0.673 0.969 [0.649] 0.951 [0.478] 0.947 [0.680] 0.969 [0.578] 0.979 [0.673]
+NQ-Open 0.726 0.639 0.925 [0.652] 0.919 [0.515] 0.939 [0.597] 0.920 [0.433] 0.976 [0.527]
+TriviaQA 0.718 0.608 0.879 [0.881] 0.863 [0.840] 0.822 [0.846] 0.735 [0.744] 0.927 [0.913]
+Table 11: Performance (Macro F1) of various evaluators across candidate LLMs and tasks. Values [in blue] represent the reference-free mode.
+E.7 CLEV in multi-reference answers
+CLEV explicitly accommodates multiple gold reference answers by incorporating all available references into the judge LLM’s prompt during evaluation.
+For datasets like AmbigQA and TriviaQA, where questions often have multiple valid answers (e.g., synonyms, rephrased answers, or alternative factual representations), CLEV aggregates all reference answers into the judge’s input prompt (e.g., concatenating them as a comma-separated list).
 This design ensures that the judge evaluates the candidate’s output against the full spectrum of acceptable answers, mirroring the human evaluation protocol, where annotators are instructed to mark a response as correct if it aligns with any reference answer.
@@
 This confusion is particularly evident in TriviaQA, where multiple reference answers introduce difficulties for the judges to recognize and evaluate a range of correct responses.
-C.9 Analysis of automatic metrics
-Figures 8, 9, 10, and 11 illustrate the fundamental trade-offs in automatic metrics.
+E.8 Analysis of automatic metrics
+Figures 3, 4, 5, and 6 illustrate the fundamental trade-offs in automatic metrics.
 In TriviaQA, where multiple normalized reference answers exist, EM achieves impressive true positives (61.7-74.3%) compared to HotpotQA (23.0-34.3%) which contains single reference answers.
@@
 For instance, despite their different contextual meanings, EM wrongly marks a match between a model prediction of “1944” (describing the start of a war) and a reference answer containing “1944” (representing the end of the war).
+Figure 3: Confusion matrices comparing the performance of automatic metrics (EM, BERTScore) and individual LLM judges on AmbigQA.
+Figure 4: Confusion matrices comparing the performance of automatic metrics (EM, BERTScore) and individual LLM judges on HotpotQA.
+Figure 5: Confusion matrices comparing the performance of automatic metrics (EM, BERTScore) and individual LLM judges on NQ-Open.
+Figure 6: Confusion matrices comparing the performance of automatic metrics (EM, BERTScore) and individual LLM judges on TriviaQA.
 EM string-matching guarantees high precision and makes EM particularly effective when exact wording is crucial, such as mathematical problems.
@@
 Additionally, false positives emerge due to the verbose responses where additional content artificially increases similarity scores.
-Figure 8: Confusion matrices comparing the performance of automatic metrics (EM, BERTScore) and individual LLM judges on AmbigQA.
-Figure 9: Confusion matrices comparing the performance of automatic metrics (EM, BERTScore) and individual LLM judges on HotpotQA.
-Figure 10: Confusion matrices comparing the performance of automatic metrics (EM, BERTScore) and individual LLM judges on NQ-Open.
-Figure 11: Confusion matrices comparing the performance of automatic metrics (EM, BERTScore) and individual LLM judges on TriviaQA.
-Appendix D Prompting
-In our main experiment, we performed zero-shot prompting in the following two stages.
-D.1 Prompting Candidate LLMs
-We prompted candidate LLMs (see Figure 12) to record generations for each task.
-We set the same role and prompt structure for each candidate model to ensure the reproducibility of our results.
-Figure 13 shows the candidate GPT-3.5-turbo response at zero temperature for the input given in Figure 12.
-You are a helpful assistant. Which comic book was also written by the writer of Crossed?
-Figure 12: Prompting candidate GPT-3.5-turbo to elicit outputs for HotpotQA.
-Crossed was written by Garth Ennis. The Boys is a comic book series also written by Garth Ennis. So the comic book that was also written by the writer of Crossed is The Boys.
-Figure 13: Candidate GPT-3.5-turbo response to the input given in Figure 12.
-D.2 Prompting LLM Judges
-We prompted LLMs-as-judges to perform the evaluation (see Figure 14). In Figure 15, judge Llama-3.1-70B evaluating candidate GPT-3.5-turbo.
-You are a helpful assistant acting as an impartial judge.
+Appendix F Prompting
+In our main experiment, we performed zero-shot prompting in the two stages.
+First, we prompted candidate LLMs to obtain outputs for the given asks.
+Then, we instructed LLM judges to perform evaluations.
+Candidate Prompt You are a helpful assistant. Which comic book was also written by the writer of Crossed?
+Figure 7: Prompt given to candidate GPT-3.5-turbo to elicit an answer for HotpotQA.
+Candidate Response Crossed was written by Garth Ennis. The Boys is a comic book series also written by Garth Ennis. So the comic book that was also written by the writer of Crossed is The Boys.
+Figure 8: Candidate GPT-3.5-turbo response to the prompt shown in Figure 7.
+Judge Prompt You are a helpful assistant acting as an impartial judge.
 You will be given a Question and a Proposed Answer.
 Your task is to judge whether the Proposed Answer is correct by comparing it to the Reference Answer.
-If the Proposed Answer is correct, choose ’True’, otherwise choose ’False’.
+If the Proposed Answer is correct, choose ’True’, otherwise, choose ’False’.
 Provide a brief explanation for your decision.
@@
 So the comic book that was also written by the writer of Crossed is The Boys.
-So the answer is: The Boys.
 Reference Answer: the Vertigo series “Preacher” Evaluation: Provide your response in the following format: Decision: [True/False] Explanation: [Your brief explanation]
-Figure 14: Prompting judge Llama-3.1-70B for evaluating candidate GPT-3.5-turbo.
-Decision: False Explanation: The Proposed Answer incorrectly identifies "The Boys" as the comic book written by the writer of Crossed.
+Figure 9: Prompting judge Llama-3.1-70B for evaluating the candidate GPT-3.5-turbo answer.
+Judge Response Decision: False Explanation: The Proposed Answer incorrectly identifies ”The Boys” as the comic book written by the writer of Crossed.
 While it is true that Garth Ennis wrote Crossed, the Proposed Answer fails to consider other works written by Garth Ennis.
-The Reference Answer, "Preacher", is indeed another comic book series written by Garth Ennis, specifically for the Vertigo imprint.
-Figure 15: Llama-3.1-70B-Judge verdict on the candidate GPT-3.5-turbo output.
+The Reference Answer, ”Preacher”, is indeed another comic book series written by Garth Ennis, specifically for the Vertigo imprint.
+Figure 10: Llama-3.1-70B-Judge verdict on the candidate GPT-3.5-turbo answer.